### v1

# The Complexity of Defining and Separating Fixpoint Formulae in Modal Logic

Abstract.
Modal separability for modal fixpoint formulae is the problem to decide for two given modal fixpoint formulae $\varphi,\varphi^{\prime}$ whether there is a modal formula $\psi$ that separates them, in the sense that $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. We study modal separability and its special case modal definability over various classes of models, such as arbitrary models, finite models, trees, and models of bounded outdegree. Our main results are that modal separability is PSpace-complete over words, that is, models of outdegree $\leq 1$, ExpTime-complete over unrestricted and over binary models, and 2-ExpTime-complete over models of outdegree bounded by some $d\geq 3$. Interestingly, this latter case behaves fundamentally different from the other cases also in that modal logic does not enjoy the Craig interpolation property over this class. Motivated by this we study also the induced interpolant existence problem as a special case of modal separability, and show that it is coNExpTime-complete and thus harder than validity in the logic. Besides deciding separability, we also provide algorithms for the effective construction of separators. Finally, we consider in a case study the extension of modal fixpoint formulae by graded modalities and investigate separability by modal formulae and graded modal formulae.
Key words and phrases:
Modal Logic, Fixpoint Logic, Separability, Interpolation, Definability
1. Introduction
For given logics $\mathcal{L},\mathcal{L}^{+}$, the $\mathcal{L}$-separability problem for $\mathcal{L}^{+}$ is to decide given two $\mathcal{L}^{+}$-formulae $\varphi,\varphi^{\prime}$ whether there is an $\mathcal{L}$-formula $\psi$ that separates $\varphi$ and $\varphi^{\prime}$ in the sense that $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. Obviously, a separator can only exist when $\varphi$ and $\varphi^{\prime}$ are mutually exclusive, and the problem is only meaningful when $\mathcal{L}$ is less expressive than $\mathcal{L}^{+}$. Intuitively, a separator formulated in a “simpler” logic $\mathcal{L}$ explains a given inconsistency in a “complicated” logic $\mathcal{L}^{+}$. Note that, for logics $\mathcal{L}^{+}$ closed under negation, $\mathcal{L}$-separability generalizes the $\mathcal{L}$-definability problem for $\mathcal{L}^{+}$: decide whether a given $\mathcal{L}^{+}$-formula is equivalent to an $\mathcal{L}$-formula. Indeed, $\varphi\in\mathcal{L}^{+}$ is equivalent to an $\mathcal{L}$-formula iff $\varphi$ and $\neg\varphi$ are $\mathcal{L}$-separable. Since separability is more general than definability, solving it requires an even better understanding of the logics under consideration. Both separability and definability are central problems with many applications in computer science. As seminal work let us only mention definability and separability of regular word languages by languages recognized by first-order logic [PZ16, Sch65, CH91].
In this paper we study definability and separability of formulae of the modal $\mu$-calculus $\mu\mathsf{ML}$ [Pra81, Koz83] by formulae in propositional modal logic $\mathsf{ML}$. $\mu\mathsf{ML}$ is the extension of $\mathsf{ML}$ with fixpoints that encompasses virtually all specification languages such as $\mathsf{PDL}$ [FL79] and $\mathsf{LTL}$ and $\mathsf{CTL}$ [BK08]. Let us consider an example.
{exa}
Consider the following properties $P_{1},P_{2},P_{3}$ of vertex-labelled trees:
  • $P_{1}$: there is an infinite path starting in the root on which each point satisfies $a$;
  • $P_{2}$: on every path there are only finitely many points satisfying $a$;
  • $P_{3}$: on every path at most two points satisfy $a$.
The properties are expressible in $\mu\mathsf{ML}$ but not in $\mathsf{ML}$, and both $P_{1},P_{2}$ and $P_{1},P_{3}$ are mutually exclusive. The properties $P_{1},P_{3}$ are separated by the $\mathsf{ML}$-formula $\psi=a\wedge\Diamond(a\wedge\Diamond a)$ which expresses that there is a path starting with three points satisfying $a$. On the other hand, no $\mathsf{ML}$-formula separates $P_{1},P_{2}$. The intuitive reason for this is that any $\mathsf{ML}$-formula $\psi$ only sees trees up to depth $|\psi|$, and one can find two trees with properties $P_{1},P_{2}$ which nonetheless look the same up to depth $|\psi|$. A formal proof of non-existence of an $\mathsf{ML}$-separator is given in Example 3 in the paper.∎
We explore the definability and separability problems over several classes of models relevant for computer science: all models, words, trees of bounded or unbounded outdegree; as well as restrictions of all these classes to finite models. On top of analyzing the decision problems, we also address the problem of constructing definitions and separators whenever they exist. The starting point for our research is the seminal paper of Otto [Ott99], where he solves modal definability over models of bounded and unbounded outdegree. In this paper, we continue this line of work, extend it to separability, and establish a fairly complete and interesting picture. Table 1 summarizes our results. We now explain its content further.
all models words binary trees $d$-ary trees, $d\geq 3$
$\mathsf{ML}$-def. ExpTime [Ott99] PSpace ExpTime [Ott99] ExpTime [Ott99]
$\mathsf{ML}$-sep. ExpTime PSpace ExpTime 2-ExpTime
sep. constr. double exp. single exp. double exp. triple exp.
interpolant ex. always always always coNExpTime
Table 1. Overview of our results. All complexity results are completeness results.
The first line essentially repeats Otto’s results; we only add the observation that $\mathsf{ML}$-definability over words is PSpace-complete. The second line contains our results for $\mathsf{ML}$-separability. The case of words is the easiest one, both in terms of computational complexity and required arguments. Next come the cases of binary and of unrestricted trees. These two classes possess some nice structural properties which (although true for different reasons) enable a common algorithmic treatment. Finally, the cases of trees with outdegree bounded by a number $d\geq 3$ enter the stage. These trees lack the good properties essential for previous constructions, which results in significantly higher computational complexity. The hardness result for $d\geq 3$ is interesting for two reasons. First, as it is entirely standard to encode trees of higher outdegree into binary ones, one could expect the ternary (and higher) case to have the same complexity as the binary one. And second, even though there are known cases when separation is provably harder than definability (regularity of visibly pushdown languages is decidable [LS19, Theorem 19] but regular separability thereof is not [Kop16, Theorem 2.4]), to the best of our knowledge our results are the only such case known in the area of logic.
The complexity landscape for deciding separability is also reflected in the maximal sizes of the separators that we construct. Relying on the well-known connection of $\mu\mathsf{ML}$ to automata, we provide effective constructions for the cases of all models, words, and binary trees. It is worth mentioning that equally effective constructions for definability over all models are given in [LQ15], but they do not work for separability. The ternary case follows from a general argument. Our construction of separators over words is optimal. Under mild assumptions (there are at least two modalities) the constructions over binary and over unrestricted trees are optimal as well, but we leave it open whether these assumptions are needed for the lower bounds. In the case of ternary and higher outdegree trees we only conjecture optimality of the constructed separators.
Finally, we observe that $\mathsf{ML}$ lacks the Craig interpolation property over trees of outdegree bounded by $d\geq 3$. Recall that a Craig interpolant for $\varphi\models\varphi^{\prime}$ in some logic $\mathcal{L}$ is a formula $\psi\in\mathcal{L}$ only using the common symbols of $\varphi$ and $\varphi^{\prime}$ and such that $\varphi\models\psi\models\varphi^{\prime}$. A logic satisfies the Craig interpolation property (CIP) if a Craig interpolant of $\varphi\models\varphi^{\prime}$ always exists. It is known that $\mathsf{ML}$ enjoys CIP over all models and over words [Gab72] and it follows from our techniques that this transfers to binary trees. In contrast and as mentioned above, over ternary and higher-arity trees $\mathsf{ML}$ lacks the CIP. It is worth mentioning that modal logic over frames of arity bounded by some $d$ has been studied under the name $\mathbf{K}\oplus\textit{{alt}}_{d}$ [Bel88]. Our results imply that $\mathbf{K}\oplus\textit{{alt}}_{d}$ enjoys CIP iff $d\leq 2$. Motivated by the lack of CIP over higher-arity trees, we study the induced interpolant existence problem – determining whether two given $\mathsf{ML}$-formulae $\varphi,\varphi^{\prime}$ admit a Craig interpolant – as a special case of separability. We show it to be coNExpTime-complete over higher arity trees, and thus harder than validity. Interpolant existence has recently been studied for other logics without CIP and our proofs are inspired by the techniques used there [JW21, AJM+23].
As an application of our results for $d$-ary trees with $d\geq 3$ we additionally present a case study: definability and separability in the graded setting in which we allow modalities saying “there are at least $k$ children such that […]” [Fin72]. Such graded modalities are a standard extension of modal logic that is especially relevant in applications in knowledge representation for conceptual modeling [BHLS17]. Hence, the larger logic $\mathcal{L}^{+}$ is the graded $\mu$-calculus in this case, see for example [KSV02, CGL99]. For the separating language $\mathcal{L}$, we consider two options. In the first, we allow graded modalities also in the separating language graded modal logic $\mathsf{grML}$, whereas in the second we stick to $\mathsf{ML}$ as before. The first setting behaves essentially as the non-graded case in the sense that both $\mathsf{grML}$-definability and $\mathsf{grML}$-separability are ExpTime-complete. More interesting is the second option, where definability and separability again show different behaviour. While $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae is ExpTime-complete, $\mathsf{ML}$-separability is 2-ExpTime-complete. The intuitive reason for the higher complexity in the latter case is that trees of bounded arity are definable in graded $\mu\mathsf{ML}$. These results are also related with a recent study about separating logics supporting counting quantifiers by logics without these [KTWZ25].
It is worth to mention that $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae generalizes the boundedness problem which asks whether a formula with a single fixpoint is equivalent to a modal formula. Boundedness has been studied for other logics such as monadic-second order logic [BOW14], datalog [HKMV95], and the guarded fragment of first-order logic [BtCCV15]. Our paper is an extension of the preliminary papers [JK24, JK25].
The paper is organized as follows. After this introduction, we set notation and recall basic facts in the preliminary Section 2. Next, we formally introduce the main notions of definability and separability, discuss a relevant construction of Otto, and solve the case of all models in Section 3. In the following Sections 4 and 5 we deal with unary and binary trees, and in Section 6 we solve the most challenging case of trees of outdegree bounded by $d\geq 3$. Section 7 applies our results to the case with graded modalities. The last Section 8 contains conclusions and final remarks.
2. Preliminaries
We recall the main notions about modal logic $\mathsf{ML}$ and the modal $\mu$-calculus $\mu\mathsf{ML}$. For the rest of this paper fix disjoint, countably infinite sets $\mathsf{Prop}$ of propositions and $\mathsf{Var}$ of variables. The syntax of $\mu\mathsf{ML}$ is given by the rule
$\varphi::=a\ |\ \neg a\ |\ \varphi\vee\varphi\ |\ \varphi\wedge\varphi\ |\ \Diamond\varphi\ |\ \Box\varphi\ |\ x\ |\ \mu x.\varphi\ |\ \nu x.\varphi$
where $a\in\mathsf{Prop}$ and $x\in\mathsf{Var}$. We assume that formulae of $\mu\mathsf{ML}$ are in a normal form such that every $x\in\mathsf{Var}$ appears at most once in a formula, and if it does appear then its appearance has a unique superformula $\psi$ beginning with $\mu x$ or $\nu x$. Modal logic $\mathsf{ML}$ is defined as the fragment of $\mu\mathsf{ML}$ with no fixpoint operators $\mu$ and $\nu$ nor variables. Both in $\mathsf{ML}$ and $\mu\mathsf{ML}$, we use abbreviations like $\top$ (for $a\vee\neg a$ for some $a\in\mathsf{Prop}$), $\Diamond^{n}\varphi$ (for a formula $\Diamond\ldots\Diamond\varphi$ with $n$ leading $\Diamond$’s), and $\neg\varphi$. A signature is a finite subset $\sigma\subseteq\mathsf{Prop}$ of propositions. We denote with $\text{sig}(\varphi)$ the set of propositions that occur in $\varphi$, and recall that the modal depth of an $\mathsf{ML}$-formula is the maximal nesting depth of $\Diamond,\Box$. With $\mathsf{ML}^{n}$ we denote the class of all $\mathsf{ML}$-formulae of modal depth at most $n$, and with $\mathsf{ML}^{n}_{\sigma}$ we denote its subclass restricted to signature $\sigma$. The set of subformulae of $\varphi$ is denoted with $\mathsf{sub}(\varphi)$. The size $|\varphi|$ of a formula $\varphi$ is the length of $\varphi$ represented as a string. This choice of the simplest possible measure of size does not matter for most of our results. We will briefly discuss alternative notions of size in the concluding Section 8.
Both $\mathsf{ML}$ and $\mu\mathsf{ML}$ are interpreted in pointed Kripke structures. More formally, a model $\mathcal{M}$ is a quadruple $\mathcal{M}=(M,v_{I},{\to},\mathsf{val})$ consisting of a set $M$ called its universe, a distinguished point $v_{I}\in M$ called the root, an accessibility relation ${\to}\subseteq M\times M$, and a valuation $\mathsf{val}:M\to\mathcal{P}(\mathsf{Prop})$.
The semantics of $\mu\mathsf{ML}$ can be defined in multiple equivalent ways. The one most convenient for us is through parity games. These are infinite-duration games with perfect information, played between players $\exists\text{ve}$ and $\forall\text{dam}$. Parity games provide a convenient definition of the $\mu\mathsf{ML}$ semantics, and will be useful in some automata-theoretic constructions. We employ a standard terminology regarding parity games, such as positions, finite and infinite plays, strategies etc. (see [Ven20] for an introduction with a special focus on $\mu\mathsf{ML}$). Additionally, given a strategy $\zeta$ for either player we call a play $\pi$ a $\zeta$-play if $\pi$ is consistent with $\zeta$. Towards the game-theoretic definition of the semantics of $\mu\mathsf{ML}$ assume a model $\mathcal{M}$ and a formula $\varphi\in\mu\mathsf{ML}$. We define a parity game called the semantic game for $\mathcal{M}$ and $\varphi$ and denoted $\mathcal{G}(\mathcal{M},\varphi)$. The positions are $M\times\mathsf{sub}(\varphi)$. The moves depend on the topmost connective. From a position of the shape $(v,\psi\vee\psi^{\prime})$ or $(v,\psi\wedge\psi^{\prime})$ it is allowed to move to either $(v,\psi)$ or $(v,\psi^{\prime})$. From $(v,\Diamond\psi)$ and $(v,\Box\psi)$ the allowed moves lead to all $(w,\psi)$ such that $v\rightarrow w$. In position $(v,a)$ or $(v,\neg a)$ the game stops and $\exists\text{ve}$ wins iff $v$ satisfies the formula component $a$ or $\neg a$, respectively. From $(v,\mu x.\psi)$ and $(v,\mu x.\psi)$ the game moves to $(v,\psi)$, and from $(v,x)$ to $(v,\psi)$ where $\psi$ is the unique superformula of $x$ beginning with $\mu x$ or $\nu x$. $\exists\text{ve}$ owns positions whose formula component has $\vee$ or $\Diamond$ as the topmost connective and $\forall\text{dam}$ owns all other positions. $\exists\text{ve}$ wins an infinite play $\pi$ if the outermost subformula seen infinitely often in $\pi$ begins with $\nu$. We say that $\mathcal{M},v$ satisfies $\varphi$ and write $\mathcal{M},v\models\varphi$ if $\exists\text{ve}$ wins the game $\mathcal{G}(\mathcal{M},\varphi)$ from position $(v,\varphi)$. Since $\mathcal{M}$ is by definition pointed, we abbreviate $\mathcal{M},v_{I}\models\varphi$ with $\mathcal{M}\models\varphi$.
The same symbol denotes entailment: $\varphi\models\psi$ means that every model of $\varphi$ is a model of $\psi$. In the case only models from some fixed class $\mathbf{C}$ are considered we talk about satisfiability and entailment over $\mathbf{C}$. Let $\mathcal{L}$ be a subset of $\mu\mathsf{ML}$ such as $\mathsf{ML}$ or $\mathsf{ML}_{\sigma}^{n}$. If two models $\mathcal{M}$ and $\mathcal{N}$ satisfy the same formulae of $\mathcal{L}$ then we call them $\mathcal{L}$-equivalent and write $\mathcal{M}\equiv_{\mathcal{L}}\mathcal{N}$.
In the paper we will study models of bounded and unbounded outdegree. The outdegree of a point $w\in M$ in a model $\mathcal{M}=(M,v_{I},\to,\mathsf{val})$ is the number of successors of $w$ in the underlying directed graph $G_{\mathcal{M}}=(M,\to)$. We say that $\mathcal{M}$ has bounded outdegree if there is a finite uniform upper bound $d$ on the outdegree of its points. In the latter case, we will call $\mathcal{M}$ $d$-ary, and binary or ternary if $d=2$ or $d=3$. If $d=1$, then we call $\mathcal{M}$ a word. A $d$-ary model is full if each of its nodes is either a leaf (i.e. has no children) or has precisely $d$ children. A model $\mathcal{M}$ is a tree if $G_{\mathcal{M}}$ is a (directed) tree with root $v_{I}$. We denote with $\mathbb{T}^{d}$ the class of all $d$-ary tree models. Both $\mathsf{ML}$ and $\mu\mathsf{ML}$ are invariant under bisimulation, and every ($d$-ary) model is bisimilar to a ($d$-ary) tree. Hence, we do not loose generality by only looking at tree models.
A prefix of a tree is a subset of its universe closed under taking ancestors. When no confusion arises we identify a prefix $N\subseteq M$ with the induced subtree $\mathcal{N}$ of $\mathcal{M}$ that has $N$ as its universe. The depth of a point is the distance from the root. The prefix of depth $n$ (or just $n$-prefix) is the set of all points at depth at most $n$ and is denoted by $M_{|_{n}}$ (and the corresponding subtree by $\mathcal{M}_{|_{n}}$).
2.1. Bisimulations
We define bisimulations and bisimilarity for trees, assuming for convenience that bisimulations only link points at the same depth. Let $\mathcal{M},\mathcal{M}^{\prime}$ be trees and $Z\subseteq M\times M^{\prime}$ a relation between $M$ and $M^{\prime}$ that relates only points of the same depth. Then, $Z$ is a bisimulation between $\mathcal{M}$ and $\mathcal{M}^{\prime}$ if it links the roots $v_{I}Zv_{I}^{\prime}$, and for every $wZw^{\prime}$ the following conditions are satisfied:
$(\mathsf{atom})$:
$\mathsf{val}(w)=\mathsf{val}^{\prime}(w^{\prime})$,
$(\mathsf{forth})$:
for every $v\in M$ with $w\to v$ there is a $v^{\prime}\in M^{\prime}$ with $w^{\prime}\to v^{\prime}$ and $vZv^{\prime}$, and
$(\mathsf{back})$:
for every $v^{\prime}\in M^{\prime}$ with $w^{\prime}\to v^{\prime}$ there is a $v\in M$ with $w\to v$ and $vZv^{\prime}$.
A functional bisimulation (also known as bounded morphism) is a function whose graph is a bisimulation. If $Z$ is a functional bisimulation from $\mathcal{M}$ to $\mathcal{M}^{\prime}$ then we write $Z:\mathcal{M}\stackrel{{\scriptstyle\mathsf{bis}}}{{\to}}{}\mathcal{M}^{\prime}$ and call $\mathcal{M}^{\prime}$ a bisimulation quotient of $\mathcal{M}$. The bisimilarity quotient of $\mathcal{M}$ is a quotient $\mathcal{M}^{\prime}$ of $\mathcal{M}$ such that if $Z^{\prime}:\mathcal{M}^{\prime}\to\mathcal{M}^{\prime\prime}$ then $\mathcal{M}^{\prime}=\mathcal{M}^{\prime\prime}$. It follows from analogous results for arbitrary models that every tree $\mathcal{M}\in\mathbb{T}^{d}$ has a unique (up to isomorphism) bisimilarity quotient $\mathcal{M}^{\prime}\in\mathbb{T}^{d}$ and that two trees are bisimilar iff their bisimilarity quotients are isomorphic.
Further, for every $n\in\mathbb{N}$ and every subset $\sigma\subseteq\mathsf{Prop}$ of the signature we consider a restricted variant of bisimulations called $(\sigma,n)$-bisimulations. In a $(\sigma,n)$-bisimulation the $\mathsf{atom}$ condition is only checked with respect to $\sigma$ and the $\mathsf{back}$ and $\mathsf{forth}$ conditions only for points at depth smaller than $n$. Formally, a relation $Z\subseteq M\times M^{\prime}$ is a $(\sigma,n)$-bisimulation if it is a bisimulation between the $n$-prefixes of the $\sigma$-reducts of $\mathcal{M},\mathcal{M}^{\prime}$. We call a $(\sigma,n)$-bisimulation between $\mathcal{M},\mathcal{M}^{\prime}$ a $(\sigma,n)$-isomorphism if it is bijective on the $n$-prefixes of $\mathcal{M},\mathcal{M}^{\prime}$. We write $\mathcal{M}\leftrightarroweq_{\sigma}^{n}\mathcal{M}^{\prime}$ if there exists a $(\sigma,n)$-bisimulation between $\mathcal{M}$ and $\mathcal{M}^{\prime}$ and $\mathcal{M}\cong_{\sigma}^{n}\mathcal{M}^{\prime}$ if there is a $(\sigma,n)$-isomorphism between them. Crucially, over every class $\mathbf{C}$ of models and for every finite $\sigma$ the equivalences $\equiv_{\mathsf{ML}_{\sigma}^{n}}$ and $\leftrightarroweq_{\sigma}^{n}$ coincide for every $n$.
2.2. Automata
We exploit the well-known connection of $\mu\mathsf{ML}$ and automata that read tree models. A nondeterministic parity tree automaton (NPTA) over $d$-ary trees $\mathbb{T}^{d}$ is a tuple $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ where $Q$ is a finite set of states, $q_{I}\in Q$ is the initial state, $\Sigma=\mathcal{P}(\sigma)$ for some signature $\sigma\subseteq\mathsf{Prop}$, $\mathsf{rank}$ assigns each state a priority, and $\delta$ is a transition function of type:
$\delta:Q\times\Sigma\to\mathcal{P}(Q^{\leq d}),$
where $Q^{\leq d}$ denotes the set of all tuples over $Q$ of length at most $d$. A run of $\mathcal{A}$ on a tree $\mathcal{M}$ is an assignment $\rho:M\to Q$ sending the root of the tree to $q_{I}$ and consistent with $\delta$ in the sense that $(\rho(v_{1}),...,\rho(v_{k}))\in\delta(\rho(v),\mathsf{val}(v)\cap\sigma)$ for every point $v$ with children $v_{1},...,v_{k}$. On occasion when considering trees of unbounded outdegree we will use automata with transition function of type
$\delta:Q\times\Sigma\to\mathcal{P}(\mathcal{P}(Q)).$
Then, consistency of $\rho$ with $\delta$ means that $\{\rho(v^{\prime})\ |\ v^{\prime}\in V\}\in\delta(\rho(v),\mathsf{val}(v)\cap\sigma)$ for every $v$ with a set $V$ of children. In either case, we call the run $\rho$ accepting if for every infinite path $v_{0},v_{1}\ldots$ in $\mathcal{M}$ the sequence $\mathsf{rank}(\rho(v_{0})),\mathsf{rank}(\rho(v_{1})),\ldots$ satisfies the parity condition. We write $\mathcal{M}\models\mathcal{A}$ in case $\mathcal{A}$ has an accepting run on $\mathcal{M}$. An automaton that is identical to $\mathcal{A}$ except that the original initial state is replaced with $q$ is denoted $\mathcal{A}[q_{I}\mapsfrom q]$. The size of an automaton $\mathcal{A}$ is the number of its states and is denoted by $|\mathcal{A}|$.
An NPTA $\mathcal{A}$ is equivalent to a formula $\varphi\in\mu\mathsf{ML}$ over a class $\mathbf{C}$ of trees when $\mathcal{M}\models\varphi$ iff $\mathcal{M}\models\mathcal{A}$ for every tree $\mathcal{M}\in\mathbf{C}$. We rely on the following classical result (see for example the discussion in [Var98] and the well-presented Dealternation Theorem 5.7 in [BC18]):
Theorem 1.
For every $\mu\mathsf{ML}$-formula $\varphi$ and class $\mathbf{C}$ of trees, we can construct an NPTA with exponentially many states equivalent to $\varphi$ over $\mathbf{C}$. The construction takes exponential time when $\mathbf{C}\subseteq\mathbb{T}^{d}$ for some $d$, and doubly exponential time in the unrestricted case.
3. Foundations of Separability
We start with recalling the notion of separability and discuss some of its basic properties. {defi} Let $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$. A formula $\psi$ separates $\varphi$ and $\varphi^{\prime}$ if $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. If $\psi$ comes from a fragment $\mathcal{L}$ of $\mu\mathsf{ML}$, then we call it an $\mathcal{L}$-separator of $\varphi$ and $\varphi^{\prime}$. It should be clear that a separator can only exist if $\varphi$ and $\varphi^{\prime}$ are inconsistent, and that either of $\varphi$ and $\neg\varphi^{\prime}$ is a trivial separator in this case. Thus, the notion is only interesting if we restrict $\mathcal{L}$. We illustrate the notion of a separator with a simple example. {exa} Consider the $\mu\mathsf{ML}$-formula $\theta_{\infty}=\nu x.\Diamond x$ expressing the existence of an infinite path originating in the root, and $\varphi^{\prime}=\Box\bot$ expressing that the root does not have any children. Clearly, $\theta_{\infty}$ and $\varphi^{\prime}$ are inconsistent, and for every $n\geq 1$ the modal formula $\Diamond^{n}\top$ separates $\theta_{\infty}$ and $\varphi^{\prime}$. Hence, there are infinitely many separators. Moreover, since $\Diamond^{m}\top\models\Diamond^{n}\top$ whenever $m\geq n$, there does not exist a logically strongest separator. ∎ The main problem we investigate in this paper is $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae, which asks whether given formulae $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ admit an $\mathsf{ML}$-separator. Since any $\mathsf{ML}$-separator of $\varphi,\neg\varphi$ is equivalent to $\varphi$, a natural special case of $\mathsf{ML}$-separability is $\mathsf{ML}$-definability, in which we assume $\varphi^{\prime}=\neg\varphi$. The notion of separation is closely related to the notion of interpolation. Let $\mathcal{L}$ be a set of $\mu\mathsf{ML}$-formulae. A Craig interpolant of $\varphi,\varphi^{\prime}\in\mathcal{L}$ is an $\mathcal{L}$-separator $\psi$ of $\varphi,\neg\varphi^{\prime}$ which only uses the common symbols of $\varphi$ and $\varphi^{\prime}$, that is, $\text{sig}(\psi)\subseteq\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$. The Craig interpolant existence problem for $\mathcal{L}$ asks if a given pair of $\mathcal{L}$-formulae admit a Craig interpolant. We say that $\mathcal{L}$ enjoys the Craig interpolation property (CIP) if every $\varphi,\varphi^{\prime}\in\mathcal{L}$ with $\varphi\models\varphi^{\prime}$ admit a Craig interpolant. In logics enjoying the CIP, the Craig interpolant existence problem to checking if $\varphi\models\varphi^{\prime}$, but in logics which lack CIP the problem may me more involved.
All introduced notions can be relativized to a class $\mathbf{C}$ of models by restricting entailment to the models in $\mathbf{C}$. In this paper, we investigate $\mathsf{ML}$-separability and $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae, as well as the Craig interpolant existence problem for $\mathsf{ML}$. The problems are considered over the class of all models and the classes $\mathbb{T}^{d}$, $d\in\mathbb{N}$ of trees of bounded outdegree.
A useful observation is that, over all classes $\mathbf{C}$ considered in this paper, if $\varphi,\varphi^{\prime}$ admit an $\mathsf{ML}$-separator $\psi$, then they admit an $\mathsf{ML}$-separator over signature $\sigma=\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. Indeed, consider $\psi_{\sigma}$ obtained from $\psi$ by replacing every non-$\sigma$ proposition with $\bot$, and $\mathbf{C}_{\sigma}$ the subclass of $\mathbf{C}$ consisting of models in which all non-$\sigma$ propositions are false everywhere. Over $\mathbf{C}_{\sigma}$, the formulae $\psi$ and $\psi_{\sigma}$ are equivalent and hence $\psi_{\sigma}$ separates $\varphi,\varphi^{\prime}$. Since every model in $\mathbf{C}$ is $\sigma$-bisimilar to a model in $\mathbf{C}_{\sigma}$, it follows that $\psi_{\sigma}$ separates $\varphi,\varphi^{\prime}$ over $\mathbf{C}$ as well. Thus, when looking for $\mathsf{ML}$-separators $\psi$ for $\varphi,\varphi^{\prime}$ we will only consider formulae over $\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. Moreover, it will be convenient to use also the notion of an $\mathsf{ML}_{\sigma}$-separator which is just a $\mathsf{ML}$-separator whose signature is contained in $\sigma$.
The starting point for the technical developments in the paper are model-theoretic characterizations for separability. Similar to what has been done in the context of interpolation, see for example [Rob60], they are given in terms of joint consistency, which we introduce next. Let $R$ be a binary relation on some class of models, such as $(\sigma,n)$-isomorphism $\cong_{\sigma}^{n}$ or $\mathsf{ML}_{\sigma}^{n}$-equivalence $\equiv_{\mathsf{ML}_{\sigma}^{n}}$. We call two formulae $\varphi,\varphi^{\prime}$ joint consistent up to $R$ (in short joint $R$-consistent) if there are models $\mathcal{M}\models\varphi$ and $\mathcal{M}^{\prime}\models\varphi^{\prime}$ with $R(\mathcal{M},\mathcal{M}^{\prime})$. For technical reasons we will sometimes also talk about joint consistency of automata $\mathcal{A},\mathcal{A}^{\prime}$ in place of formulae $\varphi,\varphi^{\prime}$. Joint $R$-consistency over a class $\mathbf{C}$ of models is defined by only looking at models from $\mathbf{C}$. Clearly, if $R^{\prime}\subseteq R$ and $\mathbf{C}^{\prime}\subseteq\mathbf{C}$ then joint $R^{\prime}$-consistency over $\mathbf{C}^{\prime}$ implies joint $R$-consistency over $\mathbf{C}$. We use the following standard equivalence:
$\varphi,\varphi^{\prime}$ are not $\mathsf{ML}_{\sigma}^{n}$-separable over $\mathbf{C}$ $\iff$ $\varphi,\varphi^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbf{C}$. ($\mathsf{Base}$)
for every $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$, $n\in\mathbb{N}$, signature $\sigma$, and class $\mathbf{C}$. The implication from right to left is immediate. The opposite one follows from the observation that for every $n\in\mathbb{N}$ and finite $\sigma$ there are only finitely many equivalence classes of $\leftrightarroweq_{\sigma}^{n}$, and each such class is fully described with a single modal formula. Let us illustrate the use of Equivalence ($\mathsf{Base}$) to show non-separability. {exa} Let $\varphi_{1}$ and $\varphi_{2}$ be $\mu\mathsf{ML}$-formulae expressing the respective properties $P_{1}$ and $P_{2}$ from Example 1. Let $\mathcal{M}$ be an infinite path in which every point satisfies $a$, and let $\mathcal{M}_{n}$ be a finite path of length $n$ in which every point satisfies $a$. Then, for each $n$ the models $\mathcal{M},\mathcal{M}_{n}$ witness joint $\leftrightarroweq^{n}_{\{a\}}$-consistency of $\varphi_{1},\varphi_{2}$. By Equivalence ($\mathsf{Base}$) this means that $\varphi_{1},\varphi_{2}$ are not $\mathsf{ML}_{\{a\}}^{n}$-separable for any $n$, and hence not $\mathsf{ML}$-separable at all. ∎
Equivalence ($\mathsf{Base}$) is also used to prove the following model-theoretic characterization of Craig interpolant existence over trees of bounded outdegree in terms of joint $\leftrightarroweq_{\sigma}^{n}$-consistency, which will be useful later on.
Lemma 2.
Let $\varphi,\varphi^{\prime}\in\mathsf{ML}^{n}$ and $\sigma=\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$. The following are equivalent for every $d\in\mathbb{N}$:
  1. (1)
    $\varphi,\varphi^{\prime}$ do not admit a Craig interpolant over $\mathbb{T}^{d}$.
  2. (2)
    $\varphi,\neg\varphi^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$.
Proof 3.1.
[…]
In the remainder of the section, we recall first, in Section 3.1, the (known) tools that were used to solve $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae, and show then, in Section 3.2 how to extend these tools (a) to solving $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae and (b) to constructing separators and definitions when they exist, both over the class of all models.
3.1. Modal Definability: A Recap
In his seminal paper [Ott99] Otto showed that $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae is ExpTime-complete over all models and over $\mathbb{T}^{d}$ for every $d\geq 2$. {thmC}[Ott99, Main Theorem and Proposition 5] Over the class of all models, as well as over $\mathbb{T}^{d}$ for every $d\geq 2$, $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae is ExpTime-complete. We start by recalling and rephrasing Otto’s construction and fixing a small mistake in the original proof. The lower bound follows by an immediate reduction from satisfiability of $\mu\mathsf{ML}$-formulae. We look at the upper bound. The first step is the following lemma, which is the heart of [Ott99, Lemma 2].
Lemma 3.
For every $\varphi\in\mu\mathsf{ML}$ and $n,d\in\mathbb{N}$ the following are equivalent:
  1. (1)
    $\varphi,\neg\varphi$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$.
  2. (2)
    $\varphi,\neg\varphi$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$.
The lemma is true, but its proof in [Ott99] is mistaken. The problem there is that the construction duplicates subtrees and hence may turn $d$-ary models into ones with outdegree greater than $d$. We present an easy alternative proof.
Proof 3.2.
[…]
Using automata-based techniques we can decide if Item 2 in Lemma 3 holds for all $n$.
Proposition 4.
Let $\mathcal{A},\mathcal{A}^{\prime}$ be parity automata and $d\in\mathbb{N}$. Then, $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n\in\mathbb{N}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{m}$-consistent over $\mathbb{T}^{d}$ for $m=|\mathcal{A}|\times|\mathcal{A}^{\prime}|+1$. The latter condition can be checked in time polynomial in $|\mathcal{A}|+|\mathcal{A}^{\prime}|$.
Proof 3.3.
(Sketch) Over $\mathbb{T}^{d}$, $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent iff they are joint $\cong_{\sigma_{0}}^{n}$-consistent for $\sigma_{0}=\sigma\cap(\text{sig}(\mathcal{A})\cup\text{sig}(\mathcal{A}^{\prime}))$. Thus, without loss of generality we assume that $\sigma\subseteq\text{sig}(\mathcal{A})=\text{sig}(\mathcal{A}^{\prime})$. Due to well-known relativization techniques we do not loose generality by only running $\mathcal{A},\mathcal{A}^{\prime}$ on full $d$-ary trees with no leaves. Let $L$ be the language of finite full $d$-ary trees over $\sigma$ such that $\mathcal{M}\in L$ iff $\mathcal{M}$ is a prefix of a $\sigma$-reduct of a model of $\mathcal{A}$. Let $L^{\prime}$ be an analogous language for $\mathcal{A}^{\prime}$. The tallness of a finite tree is the minimal distance from the root to a leaf. Observe that $\mathcal{A},\mathcal{A}^{\prime}$ are $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ iff $L\cap L^{\prime}$ contains a tree of tallness $n$. Thus, it suffices to check if $L\cap L^{\prime}$ contains trees of arbitrarily high tallness. To that end construct an automaton $\mathcal{B}$ recognizing $L\cap L^{\prime}$ of size $|\mathcal{A}|\times|\mathcal{A}^{\prime}|$. An easy pumping argument shows that the language $L\cap L^{\prime}$ of $\mathcal{B}$ contains trees of arbitrarily high tallness iff it contains a tree of tallness $m=|\mathcal{B}|+1$. To test the latter condition it is enough to inductively compute a sequence $S_{1}\supseteq S_{2}\supseteq...\supseteq S_{|\mathcal{B}|+1}$ of subsets of states of $\mathcal{B}$, where $S_{i}$ is the set of all states $q$ such that $\mathcal{B}[q_{I}\mapsfrom q]$ recognizes a tree of tallness at least $i$.
We are ready to solve $\mathsf{ML}$-definability over $\mathbb{T}^{d}$ in exponential time. Assume $\mu\mathsf{ML}$-formula $\varphi$ over signature $\sigma$. For every $n$, we know by Equivalence ($\mathsf{Base}$) that $\varphi$ is equivalent over $\mathbb{T}^{d}$ to some $\psi\in\mathsf{ML}_{\sigma}^{n}$ iff $\varphi,\neg\varphi$ are not joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$. By Lemma 3 this is equivalent to the lack of joint $\cong_{\sigma}^{n}$-consistency of $\varphi,\neg\varphi$ over $\mathbb{T}^{d}$. By Theorem 1 we can compute exponentially-sized automata $\mathcal{A}$, $\mathcal{A}^{\prime}$ equivalent to $\varphi$ and $\neg\varphi$ over $\mathbb{T}^{d}$. It follows that $\varphi$ is not $\mathsf{ML}$-definable over $\mathbb{T}^{d}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for every $n$. The last condition is decided using Proposition 4. The runtime of our algorithm is polynomial in $|\mathcal{A}|+|\mathcal{A}^{\prime}|$, and thus exponential in $|\varphi|$. This proves the part of Theorem 3.1 about $\mathbb{T}^{d}$. The remaining part concerning unrestricted models is a special case of Theorem 6, which we will prove in the following subsection.
3.2. Modal Separation: the Unrestricted Case
Over unrestricted models, separability turns out to be only slightly more complicated than definability. Lemma 3 becomes false if $\neg\varphi$ is replaced with arbitrary $\varphi^{\prime}$ (which would be the statement relevant for separability). We have the following lemma, however.
Lemma 5.
For every $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ and $n\in\mathbb{N}$ the following are equivalent:
  1. (1)
    $\varphi,\varphi^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over all models.
  2. (2)
    $\varphi,\varphi^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$, where $d=|\varphi|+|\varphi^{\prime}|$.
Proof 3.4.
[…]
$\varphi,\varphi^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over all models. (1.5)
[…]
Lemma 5 allows us to solve $\mathsf{ML}$-separability in exponential time.
Theorem 6.
Over all models, $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae is ExpTime-complete.
Proof 3.5.
The proof is almost the same as our proof of Theorem 3.1. The only difference is that we consider an arbitrary $\varphi^{\prime}$ in place of $\neg\varphi$, and hence use Lemma 5 in place of Lemma 3.
Apart from deciding separability we also construct separators when they exists. Given a subset $\mathcal{L}$ of $\mu\mathsf{ML}$-formulae, $\varphi\in\mu\mathsf{ML}$, we call $\psi\in\mathcal{L}$ an $\mathcal{L}$-uniform consequence of $\varphi$ if
$\psi\models\theta\hskip 14.22636pt\iff\hskip 14.22636pt\varphi\models\theta$
for every $\theta\in\mathcal{L}$. Note that the above condition implies $\varphi\models\psi$. The notion relativizes to a fixed class $\mathbf{C}$ of models by only considering entailment over that class. Observe that if $\varphi,\varphi^{\prime}$ are $\mathcal{L}$-separable and $\psi\in\mathcal{L}$ is an $\mathcal{L}$-uniform consequence of $\varphi$ then $\psi$ is an $\mathcal{L}$-separator for $\varphi,\varphi^{\prime}$. The same is true over any class $\mathbf{C}$.
Note that it follows from the proof of Theorem 6 that if $\varphi,\varphi^{\prime}$ are $\mathsf{ML}$-separable then they admit a separator of modal depth $n$ at most exponential in $|\varphi|+|\varphi^{\prime}|$. It follows that constructing an $\mathsf{ML}$-separator for $\varphi,\varphi^{\prime}$ boils down to constructing an $\mathsf{ML}_{\sigma}^{n}$-uniform consequence of $\varphi$ where $\sigma=\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. A naive construction which always works is to take the disjunction of all $\mathsf{ML}_{\sigma}^{n}$-types consistent with $\varphi$ over $\mathbf{C}$. Here, by an $\mathsf{ML}_{\sigma}^{n}$-type we mean a maximal consistent subset of $\mathsf{ML}_{\sigma}^{n}$. Since up to equivalence there are only finitely many formulae in $\mathsf{ML}_{\sigma}^{n}$, each $\mathsf{ML}_{\sigma}^{n}$-type can be represented as a single $\mathsf{ML}_{\sigma}^{n}$-formula and the mentioned disjunction $\psi$ is well-defined. This construction is non-elementary in $n$ over all models and doubly exponential in $n$ over models of bounded outdegree.
For some classes $\mathbf{C}$ a (i) stronger and (ii) more succinct construction is possible. Given $\varphi\in\mu\mathsf{ML}$ and $n\in\mathbb{N}$ we build $\psi\in\mathsf{ML}^{n}$ which (i) is an $\mathsf{ML}^{n}$-uniform (rather than $\mathsf{ML}_{\sigma}^{n}$-uniform) consequence of $\varphi$, and (ii) whose dependence on $n$ is only exponential. This better construction works over unrestricted models, over $\mathbb{T}^{1}$, and over $\mathbb{T}^{2}$, but notably not over $\mathbb{T}^{d}$ with $d\geq 3$. The following example shows that this gap is unavoidable, as already over $\mathbb{T}^{3}$, $\mathsf{ML}^{n}$-uniform consequences need not exist. Recall the formula $\theta_{\infty}$ from Example 3 and consider $\varphi=\Diamond(\theta_{\infty}\wedge a)\wedge\Diamond(\neg\theta_{\infty}\wedge a)$ which says that there are at least two children satisfying $a$: one with an infinite path and the other without. Over ternary models $\varphi$ enforces that there is at most one child satisfying $\neg a$, and hence entails $\theta_{c}=\Diamond(\neg a\wedge c)\Rightarrow\Box(\neg a\Rightarrow c)$ for every $c$. It is easy to check that over $\mathbb{T}^{3}$ any consequence $\psi\in\mathsf{ML}$ of $\varphi$ can only entail $\theta_{c}$ when $c\in\text{sig}(\psi)$. Consequently, $\varphi$ has no $\mathsf{ML}^{1}$-uniform consequence over $\mathbb{T}^{3}$.
We now go back to the construction of $\mathsf{ML}^{n}$-uniform consequences in the feasible cases: all models, $\mathbb{T}^{1}$, and $\mathbb{T}^{2}.$ Since in the following Section 4 we will provide a more efficient construction for $\mathbb{T}^{1}$, we focus on the unrestricted and binary case. For convenience, we construct $\mathsf{ML}^{n}$-uniform consequences of automata instead of formulae, with definition adapted in an obvious way. In the binary case we will assume that automata are duplication safe meaning that for every singleton transition $(p)\in\delta(q,c)$ its duplication is also legal $(p,p)\in\delta(q,c)$. Every automaton equivalent to a $\mu\mathsf{ML}$-formula can be made duplication safe.
Proposition 7.
Let $\mathbf{C}$ be the class of all models or $\mathbb{T}^{2}$. Assume an NPTA $\mathcal{A}$ over $\mathbf{C}$, and in the case $\mathbf{C}=\mathbb{T}^{2}$ assume that $\mathcal{A}$ is duplication safe. For every $n\in\mathbb{N}$, an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbf{C}$ can be constructed in time $|\mathcal{A}|^{O(n\cdot|\mathcal{A}|)}$ if $\mathbf{C}$ is the class of all models and in time $2^{O(n\cdot|\mathcal{A}|)}$ if $\mathbf{C}=\mathbb{T}^{2}$.
Proof 3.6.
Let $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ be an NPTA over $\mathbf{C}$. In the case $\mathbf{C}=\mathbb{T}^{2}$ assume that $\mathcal{A}$ is duplication safe. We construct $\psi_{m,q}$ for every $q\in Q$ and $m\leq n$ by induction on $m$. For the base case we put:
$\psi_{0,q}=\bigvee\{c\in\Sigma\ |\ \text{there is $\mathcal{N}\in\mathbf{C}$ with $\mathcal{N}\models\mathcal{A}[q_{I}\mapsfrom q]$ and $\mathcal{N}\models c$}\}$
For the induction step define:
$\psi_{m+1,q}=\bigvee_{c\in\Sigma}\bigvee_{S\in\delta(q,c)}c\wedge\nabla\{\psi_{m,p}\ |\ p\in S\}$
where $\nabla\Phi$ is an abbreviation for $\bigwedge_{\theta\in\Phi}\Diamond\theta\wedge\Box\bigvee_{\theta\in\Phi}\theta$. Assume $\mathbf{C}$ is either the class of all models or $\mathbb{T}^{2}$. The construction preserves the following invariant:
[…]
for every structure $\mathcal{M}\in\mathbf{C}$. Hence, $\psi_{n,q_{I}}$ is an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbf{C}$. It is routine to check that the formula $\psi_{n,q_{I}}$ has the right size.
The proof of (1) proceeds by induction. The base case is immediate so we focus on the inductive case. We start with the easier implication $\Leftarrow$ from right to left. Assume $\mathcal{M}\leftrightarroweq^{m+1}\mathcal{N}$ for some $\mathcal{N}\in\mathbf{C}$ such that $\mathcal{N}\models\mathcal{A}[q_{I}\mapsfrom q]$. Let $c$ and $N_{0}$ denote respectively the color and the set of all children of the root of $\mathcal{N}$. There is an accepting run $\rho$ of $\mathcal{A}[q_{I}\mapsfrom q]$ on $\mathcal{N}$. We have $\rho[N_{0}]=S$ for some transition $S\in\delta(q,c)$. For each $v\in N_{0}$ the run $\rho$ witnesses that the subtree of $\mathcal{N}$ rooted at $v$ is accepted by $\mathcal{A}[q_{I}\mapsfrom\rho(v)]$. By induction hypothesis, this means that the subtree satisfies $\psi_{\rho(v),m}$. It follows that $\mathcal{N}$ satisfies $\nabla\{\psi_{m,p}\ |\ p\in S\}$ and hence also $\psi_{q,m+1}$. By $\mathcal{M}\leftrightarroweq^{m+1}\mathcal{N}$ this implies $\mathcal{M}\models\psi_{q,m+1}$.
We now prove the other implication $\Rightarrow$ from left to right. The proof uses similar but slightly different arguments in the cases of binary and unrestricted models. It is worth to point out, however, that the implication $\Rightarrow$ would not be valid over $\mathbb{T}^{d}$ with $d\geq 3$. Assume $\mathcal{M}\models\psi_{m+1,q}$, let $c$ be the color of the root and $M_{0}$ its children. There is $S\in\delta(q,c)$ such that the nabla $\nabla$ of $\Phi=\{\psi_{m,p}\ |\ p\in S\}$ is satisfied in the root. This means that the subtree of each $v\in M_{0}$ satisfies some $\psi_{m,p}$, and conversely each $\psi_{m,p}$ is satisfied in some of these subtrees.
We define a model $\mathcal{N}^{\prime}$. Its exact definition depends on which class $\mathbf{C}$ of models we consider. If $\mathbf{C}$ is the class of all models we obtain $\mathcal{N}^{\prime}$ from $\mathcal{M}$ by replacing the subtree of each $v\in M_{0}$ with $|Q|$ distinct copies of that subtree. If $\mathbf{C}=\mathbb{T}^{2}$ then $\mathcal{N}^{\prime}$ is the full binary tree obtained from $\mathcal{M}$ by duplicating subtrees. In either case, let $N_{0}^{\prime}$ denote the children of the root of $\mathcal{N}^{\prime}$. By design $\mathcal{N}^{\prime}$ has a separate witness for each formula in $\Phi$. That is, there is a surjective assignment $h:N_{0}^{\prime}\to\Phi$ which maps every $v\in N_{0}^{\prime}$ to some formula $\psi_{m,p}$ true in the subtree of $v$.
[…]
Given the exponential construction of automata from Theorem 1 and the exponential upper bound on modal depth $n$ of separators, Proposition 7 yields an efficient construction of separators.
Theorem 8.
If $\varphi,\varphi^{\prime}$ are $\mathsf{ML}$-separable, then one can compute an $\mathsf{ML}$-separator in time doubly exponential in $|\varphi|+|\varphi^{\prime}|$.
It is not difficult to show that, in the presence of at least two accessibility relations $\Diamond_{1},\Diamond_{2}$, the construction is optimal: one can express in $\mu\mathsf{ML}$ that the model embeds a full binary tree of depth $2^{n}$ and in which each inner node has both a $\Diamond_{1}$- and a $\Diamond_{2}$-successor. Using standard techniques, one can show that any modal formula expressing this property is of doubly exponential size [FvdHIK13]. Whether having two accessibility relations is necessary for this lower bound is an interesting question which we leave open.
The remaining open cases are the problems of $\mathsf{ML}$-separability (and separator construction) over $\mathbb{T}^{d}$ for $d\geq 1$. We investigate the cases of unary ($d=1$), binary ($d=2$), and higher maximal outdegree ($d\geq 3$) in turn, starting in Section 4. We emphasize that the outdegree $d$ is not a part of the input but rather a property of the considered class of models.
4. Unary Case
We first investigate $\mathsf{ML}$-separability over $\mathbb{T}^{1}$, that is, models that are essentially finite or infinite words. Note that satisfiability of $\mu\mathsf{ML}$ over words is PSpace-complete: an upper bound follows, e.g. via the translation to automata and the lower bound is inherited from $\mathsf{LTL}$ [SC85, Theorem 4.1]. This suggests that also definability and separability could be easier. Indeed, we show:
Theorem 9.
$\mathsf{ML}$-definability and $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae are PSpace-complete over $\mathbb{T}^{1}$.
Proof 4.1.
The lower bound is by a reduction from satisfiability, and applies to definability. Indeed, let $\varphi\in\mu\mathsf{ML}$ and let $\psi=\mu x.(a\vee\Diamond x)$ be the $\mu\mathsf{ML}$-formula expressing that $a$ is satisfied at some point in the model, for a proposition $a$ not occurring in $\varphi$. It is easy to see that $\varphi$ is unsatisfiable over $\mathbb{T}^{1}$ iff $\varphi\wedge\psi$ is $\mathsf{ML}$-definable (namely $\bot$).
For the upper bound, consider formulae $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ and let $\sigma=\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. Define the set $L$ of finite words over $\Sigma=\mathcal{P}(\sigma)$ by taking
[…]
and let $L^{\prime}$ be analogously defined for $\varphi^{\prime}$. Note that two models in $\mathbb{T}^{1}$ are bisimilar iff they are isomorphic. Hence, by Equivalence ($\mathsf{Base}$), $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ are not $\mathsf{ML}$-separable over $\mathbb{T}^{1}$ iff $L\cap L^{\prime}$ is infinite. It is standard to define a finite automaton $\mathcal{A}$ recognizing $L\cap L^{\prime}$ and check if its language is infinite (which is equivalent to checking if $L\cap L^{\prime}$ contains input longer than $|\mathcal{A}|$). To do it in polynomial space, we nondeterministically guess the long input, letter by letter, and only remember the current state and a binary counter measuring the length of the input guessed so far.
We conclude this section with proving that $\mathsf{ML}$-separators can be constructed in exponential time and are thus of at most exponential size. Note that this is optimal, since over $\mathbb{T}^{1}$, $\mu\mathsf{ML}$ is exponentially more succinct than $\mathsf{ML}$. Indeed, it is standard to implement an exponential counter using a polynomially sized $\mu\mathsf{ML}$-formula.
Theorem 10.
If $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ are $\mathsf{ML}$-separable over $\mathbb{T}^{1}$, then one can compute an $\mathsf{ML}$-separator in time exponential in $|\varphi|+|\varphi^{\prime}|$.
As argued in the previous section, it suffices to construct an $\mathsf{ML}^{n}$-uniform consequence of the NPTA equivalent to $\varphi$, which we do next. {restatable}proppropluniformone Let $\mathcal{A}$ be an NPTA over $\mathbb{T}^{1}$. For every $n\in\mathbb{N}$, an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbb{T}^{1}$ can be constructed in time polynomial in $n$ and $|\mathcal{A}|$.
Proof 4.2.
Take an NPTA $\mathcal{A}=(Q,\Sigma,\delta,q_{I},\mathsf{rank})$ and $n\in\mathbb{N}$. We first construct, for every $m\leq n$ and $p,q\in Q$, formulae $\psi^{m}_{pq}\in\mathsf{ML}_{\sigma}^{n}$, such that for every $\mathcal{M}\in\mathbb{T}^{1}$:
[…]
The definition is by induction on $m$:
[…]
$\psi^{1}_{pq}$ $=\bigvee\{c\ \mid\ c\in\Sigma,\{q\}\in\delta(p,c)\}$
$\psi^{m}_{pq}$ $=\bigvee_{q^{\prime}\in Q}\big(\psi^{\lfloor m/2\rfloor}_{pq^{\prime}}\wedge\Diamond^{\lfloor m/2\rfloor}\psi^{\lceil m/2\rceil}_{q^{\prime}q}\big)\quad\text{for $1<m\leq n$}$
It is routine to verify that $\psi^{m}_{pq}$ satisfies (2) and is of size $|\psi^{m}_{pq}|\in O(|Q|\cdot m^{2})$.
In the construction of the $n$-uniform consequence of $\mathcal{A}$, a bit care has to be taken since $\mathcal{A}$ may accept words shorter or longer than $n$. To deal with this, we introduce some more notation. Since we are working over $\mathbb{T}^{1}$, $\delta(q,c)$ contains only sets of cardinality at most $1$. The case $\emptyset\in\delta(q,c)$ is of particular interest because this means that the automaton can accept in state $q$ reading color $c$. Denote with $\textit{Acc}_{q}$ the set of $c\in\Sigma$ with $\emptyset\in\delta(q,c)$. Further denote with $\textit{Cont}_{q}$ the set of all $c$ such that $\mathcal{A}[q_{I}\mapsfrom q]$ accepts a word starting with $c$. We finish the construction by setting:
$\psi_{n}=\bigvee_{q\in Q}\left(\psi^{n}_{q_{I}q}\wedge\Box^{n}\bigvee_{c\in\textit{Cont}_{q}}c\right)\vee\bigvee_{m\leq n}\bigvee_{q\in Q}\left(\psi^{m}_{q_{I}q}\wedge\Box^{m+1}\bot\wedge\Box^{m}\bigvee_{c\in\textit{Acc}_{q}}c\right).$
Intuitively, $\psi_{n}$ describes all possible prefixes of length $n$ of words accepted by $\mathcal{A}$. It is readily checked that $\psi_{n}$ satisfies the required size bounds and that $\mathcal{A}\models\psi_{n}$. To verify that $\psi_{n}\models\theta$ for every $\theta\in\mathsf{ML}^{n}$ with $\mathcal{A}\models\theta$, we show the following equivalence for all $\mathcal{M}\in\mathbb{T}^{1}$:
$\mathcal{M}\models\psi_{n}\text{\hskip 14.22636pt $\iff$ \hskip 14.22636pt there exists $\mathcal{N}\models\mathcal{A}$ with $\mathcal{N}\leftrightarroweq^{n}\mathcal{M}$.}$ (3)
For $\Rightarrow$”, fix $\mathcal{M}\in\mathbb{T}^{1}$ with $\mathcal{M}\models\psi_{n}$. If $\mathcal{M}\models\psi^{n}_{q_{I}q}\wedge\Box^{n}\bigvee_{c\in\textit{Cont}_{q}}c$ for some $q$, then by Invariant (2), there is a run of $\mathcal{A}$ from the initial state $q_{0}$ to some state $q\in Q$ when reading the $n$-prefix of $\mathcal{M}$, and the last color in the prefix is $c$. Since $c\in\textit{Cont}_{q}$, we can extend the $n$-prefix of $\mathcal{M}$ to an $\mathcal{N}\in\mathbb{T}^{1}$ accepted by $\mathcal{A}$. Clearly, in this way $\mathcal{N}\leftrightarroweq^{n}\mathcal{M}$. If $\mathcal{M}\models\psi^{m}_{q_{I}q}\wedge\Box^{m+1}\bot\wedge\Box^{m}\bigvee_{c\in\textit{Acc}_{q}}c$, for some $m\leq n$ and $q\in Q$, then $\mathcal{M}$ is a finite word of length $m$ that is accepted by the automaton. We can take $\mathcal{N}=\mathcal{M}$ in this case.
For $\Leftarrow$”, let $\mathcal{M}\in\mathbb{T}^{1}$ a word such that there is some $\mathcal{N}\models\mathcal{A}$ with $\mathcal{N}\leftrightarroweq^{n}\mathcal{M}$. The former condition $\mathcal{N}\models\mathcal{A}$ implies the existence of an accepting run $\rho$ of $\mathcal{A}$ on $\mathcal{N}$. The latter, $\mathcal{N}\leftrightarroweq^{n}\mathcal{M}$, implies that $\mathcal{N}$ and $\mathcal{M}$ coincide on their $n$-prefixes. We distinguish cases. If the length of $\mathcal{N}$ is greater than $n$, then the $n$-prefix of $\rho$ ending in state $q$ witnesses $\mathcal{M}\models\psi^{n}_{q_{I}q}\wedge\Box^{n}\bigvee_{c\in\textit{Cont}_{q}}c$. Otherwise, the depth of $\mathcal{N}$ is $m\leq n$ and the run $\rho$ ending in $q$ witnesses that $\mathcal{M}\models\psi^{m}_{q_{I}q}\wedge\Box^{m+1}\bot\wedge\Box^{m}\bigvee_{c\in\textit{Acc}_{q}}c$.
5. Binary Case
We next handle the binary case $\mathbb{T}^{2}$. The key observation here is that, between full binary trees, bisimilarity entails isomorphism.
Proposition 11.
Assume full binary trees $\mathcal{M},\mathcal{M}^{\prime}\in\mathbb{T}^{2}$. If $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are $\sigma$-bisimilar then they are $\sigma$-isomorphic.
Proof 5.1.
By definition a $\sigma$-bisimulation between two models is a bisimulation between their reducts to $\sigma$, and $\sigma$-isomorphism is such a bisimulation which is additionally bijective. It therefore suffices to show that if $\mathcal{M},\mathcal{M}^{\prime}$ are full binary trees and $Z$ is a bisimulation between them then there is a bijective bisimulation $Z^{\prime}\subseteq Z$. We construct such a $Z^{\prime}$. To that end, we inductively construct a descending sequence:
$Z\supseteq Z_{1}\supseteq Z_{2}\supseteq Z_{3}\supseteq$...
of bisimulations such that for each $n$ the restriction of $Z_{n}$ to the $n$-prefixes of $\mathcal{M}$ and $\mathcal{M}^{\prime}$ is bijective. The induction base $n=1$ is trivial with $Z_{1}=Z$. For the induction step $n+1$ let $Z_{n}\subseteq Z$ be the bisimulation given by the inductive hypothesis. $Z_{n}$ bijectively maps the points $v_{1},...,v_{k}$ at depth $n$ in $\mathcal{M}$ to the respective points $w_{1},...,w_{k}$ at depth $n$ in $\mathcal{M}^{\prime}$. For each $i$ we have $v_{i}Zw_{i}$. Hence, either both $v_{i}$ and $w_{i}$ are leaves (i), or both have respective children $v_{i}^{l},v_{i}^{r}$ and $w_{i}^{l},w_{i}^{r}$. In the latter case either (ii) $v_{i}^{l}Zw_{i}^{l}$ and $v_{i}^{r}Zw_{i}^{r}$ or (iii) $v_{i}^{l}Zw_{i}^{r}$ and $v_{i}^{r}Zw_{i}^{l}$ (the cases (ii) and (iii) are not exclusive). Consider the bijective relation $K^{i}\subseteq Z$ between children of $v_{i}$ and children of $w_{i}$:
$K^{i}=\begin{cases}\emptyset&\text{if $v_{i}$ and $w_{i}$ are leaves (i),}\\ \{(v_{i}^{l},w_{i}^{l}),(v_{i}^{r},w_{i}^{r})\}&\text{if (ii),}\\ \{(v_{i}^{l},w_{i}^{r}),(v_{i}^{r},w_{i}^{l})\}&\text{otherwise.}\end{cases}$
The bisimulation $Z_{n+1}$ is constructed as follows. It is identical to $Z_{n}$ between points at levels at most $n$, to $\bigcup_{i\leq k}K^{i}$ between points at level exactly $n+1$, and to $Z$ between points at strictly greater levels. No points at mismatching levels are linked. It is straightforward to verify that such $Z_{n+1}\subseteq Z_{n}$ is a bisimulation, and that its restriction to the $n+1$-prefixes of $\mathcal{M}$ and $\mathcal{M}^{\prime}$ is bijective.
We conclude the proof by taking the limit $Z^{\prime}=\bigcap_{n\in\mathbb{N}}Z_{n}$ as the desired bijective bisimulation between $\mathcal{M}$ and $\mathcal{M}^{\prime}$.
Proposition 11 can be used to prove the Craig interpolation property of $\mathsf{ML}$ over $\mathbb{T}^{2}$.
Proposition 12.
$\mathsf{ML}$ over $\mathbb{T}^{2}$ has the Craig interpolation property.
Proof 5.2.
Assume $\varphi,\varphi^{\prime}\in\mathsf{ML}$, let $n$ be the maximum of their modal depths and denote $\sigma=\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$. Assume towards contradiction that $\varphi\models\varphi^{\prime}$ but $\varphi,\varphi^{\prime}$ do not have a Craig interpolant. By Lemma 2 the latter implies that $\varphi,\neg\varphi^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{2}$. Let $\mathcal{M}\leftrightarroweq_{\sigma}^{n}\mathcal{M}^{\prime}$ be respective binary trees witnessing this. By duplicating subtrees and trimming if needed, we may assume that $\mathcal{M},\mathcal{M}^{\prime}$ are full and have depth at most $n$, which therefore by Proposition 11 implies $\mathcal{M}\cong_{\sigma}\mathcal{M}^{\prime}$. Let $\mathcal{N}$ be a model obtained from $\mathcal{M},\mathcal{M}^{\prime}$ by taking their common underlying graph, and valuation of every proposition $a\in\text{sig}(\varphi)$ or $a\in\text{sig}(\varphi^{\prime})$ inherited from $\mathcal{M}$ or $\mathcal{M}^{\prime}$, respectively (by $\mathcal{M}\cong_{\sigma}\mathcal{M}^{\prime}$ this is well-defined for $a\in\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$). It follows that $\mathcal{M}\cong_{\text{sig}(\varphi)}\mathcal{N}\cong_{\text{sig}(\varphi^{\prime})}\mathcal{M}^{\prime}$ and hence $\mathcal{N}$ satisfies both $\varphi$ and $\neg\varphi^{\prime}$, a contradiction.
Apart from the above Proposition 12, Proposition 11 also implies the following separability-variant of Lemma 3 over $\mathbb{T}^{2}$.
Lemma 13.
For every $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ and $n\in\mathbb{N}$ the following are equivalent:
  1. (1)
    $\varphi,\varphi^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{2}$.
  2. (2)
    $\varphi,\varphi^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{2}$.
Proof 5.3.
We show only the nontrivial implication 1 $\Rightarrow$ 2. Assume binary $\mathcal{M}\models\varphi$, $\mathcal{M}^{\prime}\models\varphi^{\prime}$ with $\mathcal{M}\leftrightarroweq_{\sigma}^{n}\mathcal{M}^{\prime}$. Let $\mathcal{N}\models\varphi$ and $\mathcal{N}^{\prime}\models\varphi^{\prime}$ be full binary trees obtained from $\mathcal{M}$ and $\mathcal{M}^{\prime}$ by duplicating subtrees. By Proposition 11, $\mathcal{N}\cong_{\sigma}^{n}\mathcal{N}^{\prime}$ which proves 2.
Similarly to the definability case, Lemma 13 combined with Equivalence ($\mathsf{Base}$) and Proposition 4 immediately give an exponential procedure for separability. Since the lower bound is inherited from definability, we get the following result.
Theorem 14.
$\mathsf{ML}$-separability and $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae are ExpTime-complete over $\mathbb{T}^{2}$.
With the same argument as for Theorem 8 we use Proposition 7 to conclude:
Theorem 15.
If $\varphi,\varphi^{\prime}$ are $\mathsf{ML}$-separable over $\mathbb{T}^{2}$, then one can compute an $\mathsf{ML}$-separator in time doubly exponential in $|\varphi|+|\varphi^{\prime}|$.
6. Ternary and Beyond
In this section we address the case of models with outdegree bounded by a number $d\geq 3$. We illustrate that this case behaves differently as it lacks the Craig interpolation property.
Theorem 16.
For $d\geq 3$, $\mathsf{ML}$ over $\mathbb{T}^{d}$ does not enjoy the Craig interpolation property.
Proof 6.1.
We give the proof for $d=3$, the proof for $d>3$ is similar. Consider $\mathsf{ML}$-formulae $\varphi=\Diamond(a\wedge b)\wedge\Diamond(a\wedge\neg b)$ and $\varphi^{\prime}=\Diamond(\neg a\wedge c)\wedge\Diamond(\neg a\wedge\neg c)$. Clearly, $\varphi\models\neg\varphi^{\prime}$ over $\mathbb{T}^{3}$. Observe that models $\mathcal{M},\mathcal{M}^{\prime}$ in Figure 1 witness that $\varphi,\varphi^{\prime}$ are joint $\leftrightarroweq_{\{a\}}$-consistent and thus joint $\leftrightarroweq_{\{a\}}^{n}$-consistent for every $n\in\mathbb{N}$. By Equivalence ($\mathsf{Base}$) there is no $\mathsf{ML}_{\{a\}}$-separator, which is nothing else than a Craig interpolant.
$\mathcal{M}\mathcal{M}^{\prime}$$v_{I}$$ba,\neg b¬a\neg a$$v_{I}^{\prime}$$\neg a,c$$\neg a,\neg c$$a$
Figure 1. Witness of joint consistency: dashed lines and colors indicate the $\{a\}$-bisimulation.
Motivated by the lack of the Craig interpolation property, we study the Craig interpolant existence problem for $\mathsf{ML}$ over $\mathbb{T}^{d}$ with $d\geq 3$. We show that this problem is harder than $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae over arbitrary models.
{restatable}
thmthmmlinterpolants For $d\geq 3$, the Craig interpolant existence for $\mathsf{ML}$ over $\mathbb{T}^{d}$ is coNExpTime-complete. We start with the upper bound, for any fixed $d\geq 3$. Let $\varphi,\varphi^{\prime}\in\mathsf{ML}^{n}$ and let $\sigma=\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$ be the common signature. By Lemma 2 and using the observation that witnesses for joint $\leftrightarroweq_{\sigma}^{n}$-consistency over $\mathbb{T}^{d}$ can be assumed to be of depth at most $n$, the non-existence of a Craig interpolant for $\varphi,\neg\varphi^{\prime}$ can be decided by a standard “guess-and-check”-procedure:
  1. (1)
    Guess two structures $\mathcal{M},\mathcal{M}^{\prime}\in\mathbb{T}^{d}$ of depth $n$ in which no point satisfies a non-$\sigma$ propositions;
  2. (2)
    Verify that $\mathcal{M}\models\varphi$, $\mathcal{M}^{\prime}\models\neg\varphi^{\prime}$, and $\mathcal{M}\leftrightarroweq^{n}_{\sigma}M^{\prime}$.
The runtime of the procedure is exponential: any structure in (1) can be represented in size polynomial in $|\sigma|$ and $d^{n}$. Moreover, model checking in modal logic and bisimulation testing is possible in polynomial time in the size of the structure and the given formulae.
The lower bound is more intriguing and relies on an extension of the counterexample to the Craig interpolation property in the proof of Theorem 16. For the sake of simplicity, we do the proof for $d=3$; the proof for $d>3$ is similar. Reconsidering the proof of Theorem 16 it is important to note that in every witness $\mathcal{M},\mathcal{M}^{\prime}$ of joint $\leftrightarroweq_{\{a\}}^{n}$-consistency of $\varphi,\varphi^{\prime}$, there are two successors of $v_{I}$ that are bisimilar to the same successor of $v_{I}^{\prime}$. We extend the idea and enforce exponentially many bisimilar points by providing suitable families of formulae $(\psi_{i})_{i\in\mathbb{N}},(\psi_{i}^{\prime})_{i\in\mathbb{N}}$.
Lemma 17.
There are families of formulae $(\psi_{i})_{i\in\mathbb{N}},(\psi_{i}^{\prime})_{i\in\mathbb{N}}$ such that:
  1. (1)
    The size of the formulae $\psi_{i},\psi_{i}^{\prime}$ is polynomial in $i$.
  2. (2)
    $\text{sig}(\psi_{i})=\{a,b_{0},\ldots,b_{i-1}\}$ and $\text{sig}(\psi_{i}^{\prime})=\{a,c\}$.
  3. (3)
    For every $i\in\mathbb{N}$, for every $\mathcal{M},\mathcal{M}^{\prime}\in\mathbb{T}^{3}$ with $\mathcal{M}\models\psi_{i}$, $\mathcal{M}\models\psi_{i}^{\prime}$, for every signature $\tau$ with $a\in\tau$ and $\{c,b_{0},\ldots,b_{i-1}\}\cap\tau=\emptyset$, and every $(\tau,\ell)$-bisimulation $Z$ witnessing $\mathcal{M}\leftrightarroweq^{\ell}_{\tau}\mathcal{M}^{\prime}$ for some $\ell\geq i$, there are points $w_{0},\ldots,w_{2^{i}-1}$ in depth $i$ in $\mathcal{M}$ and a point $\widehat{w}$ in depth $i$ in $\mathcal{M}^{\prime}$ such that:
    1. (a)
      $(w_{j},\widehat{w})\in Z$ for all $j$,
    2. (b)
      all $w_{j}$ and $\widehat{w}$ satisfy $a$, and
    3. (c)
      distinct $w_{j},w_{k}$ can be distinguished by some proposition in $b_{0},\ldots,b_{i-1}$.
Proof 6.2.
Define the families $(\psi_{i})_{i\in\mathbb{N}}$, $(\psi_{i}^{\prime})_{i\in\mathbb{N}}$ of modal formulae inductively as follows:
$\psi_{0}$ $=\psi_{0}^{\prime}=a$
$\psi_{i+1}$ $=\Diamond(a\wedge b_{i})\wedge\Diamond(a\wedge\neg b_{i})\wedge\Box\big(a\to(\psi_{i}\wedge(b_{i}\to\textstyle\bigwedge_{j<i}\Box^{j}b_{i})\wedge(\neg b_{i}\to\textstyle\bigwedge_{j<i}\Box^{j}\neg b_{i}))\big)$
$\psi_{i+1}^{\prime}$ $=\Diamond(\neg a\wedge\textstyle\bigwedge_{j<i}\Box^{j}\neg a\wedge c)\wedge\Diamond(\neg a\wedge\bigwedge_{j<i}\Box^{j}\neg a\wedge\neg c)\wedge\Diamond(a\wedge\psi_{i}^{\prime})$
For Item (1), it is routine to verify that both $\psi_{i}$ and $\psi_{i}^{\prime}$ are of quadratic size. Item (2) is immediate from the definition. Item (3) is shown by induction on $i$, using the same arguments from the proof of Theorem 16. The base case $i=0$ is trivial. For the induction step, let $\mathcal{M},\mathcal{M}^{\prime}\in\mathbb{T}^{3}$ with $\mathcal{M}\models\psi_{i+1}$, $\mathcal{M}^{\prime}\models\psi_{i+1}^{\prime}$, let $\tau$ be any signature with $a\in\tau$ and $\{c,b_{0},\ldots,b_{i-1}\}\cap\tau=\emptyset$, and let $Z$ be a $(\tau,\ell)$-bisimulation witnessing $\mathcal{M}\leftrightarroweq^{\ell}_{\tau}\mathcal{M}^{\prime}$ for some $\ell\geq i+1$. By the definition of $\psi_{i+1}$, there are distinct successors $v_{1},v_{2}$ of the root $v_{I}$ of $\mathcal{M}$ both satisfying $\psi_{i}$. By the definition of $\psi_{i+1}^{\prime}$, there is a successor $v^{\prime}$ of the root $v_{I}^{\prime}$ of $\mathcal{M}^{\prime}$ satisfying $\psi_{i}^{\prime}$. By the same argument as in the proof of Theorem 16, we must have $v_{1}Zv^{\prime}$ and $v_{2}Zv^{\prime}$. Applying the induction hypothesis to $v_{1}$ and $v^{\prime}$ gives us points $w_{0},\ldots,w_{2^{i}-1}$ in depth $i$ below $v_{1}$ in $\mathcal{M}$ and a point $\widehat{w}$ in depth $i$ below $v^{\prime}$ in $\mathcal{M}^{\prime}$ satisfying 3(a)–(c). Applying the induction hypothesis to $v_{2}$ and $v^{\prime}$ gives us points $w_{0}^{\prime},\ldots,w_{2^{i}-1}^{\prime}$ in depth $i$ below $v_{2}$ in $\mathcal{M}^{\prime}$ and a point $\widehat{w}^{\prime}$ in depth $i$ below $v^{\prime}$ in $\mathcal{M}^{\prime}$ satisfying 3(a)–(c). Note that $\widehat{w}=\widehat{w}^{\prime}$ since, by 3(b), both satisfy $a$ and the definition of $\psi_{i+1}^{\prime}$ ensures that there is a unique element satisfying $a$ at every depth below $v^{\prime}$. The sequence $w_{0},\ldots,w_{2^{i}-1},w_{0}^{\prime},\ldots,w_{2^{i}-1}^{\prime}$ together with $\widehat{w}$ satisfy 3(a)–(c) for $i+1$. This concludes the induction step and thus the proof of Item (3).
Intuitively, this means that $\psi_{i},\psi_{i}^{\prime}$ from Lemma 17 enforce in joint $\leftrightarroweq_{\{a\}}^{i}$-consistent models $\mathcal{M},\mathcal{M}^{\prime}$ that $\mathcal{M}$ contains $2^{i}$ points $w_{0},\ldots,w_{2^{i}-1}$ which are all linked to the same point $\widehat{w}$ in $\mathcal{M}^{\prime}$. We exploit this link to synchronize information between the $w_{j}$, following a strategy that has recently been used to show coNExpTime-hardness for interpolant existence in some description logics [AJM+23].
We reduce a suitable torus tiling problem. Let $\Delta$ be a (finite) set of tile types, $H,V\subseteq\Delta\times\Delta$ be horizontal and vertical compatibility relations, and let $n\in\mathbb{N}$. Then a function $\tau:\{0,\ldots,n-1\}\times\{0,\ldots,n-1\}\to\Delta$ is a torus tiling for $\Delta,H,V,n$ if for all $x,y\in\{0,\ldots,n-1\}$:
  • $(\tau(x,y),\tau(x,y\oplus_{n}1))\in H$, and
  • $(\tau(x,y),\tau(x\oplus_{n}1,y))\in V$,
where $\oplus_{n}$ denotes addition modulo $n$. It is well-known that the exponential torus tiling problem, which takes as input $\Delta,H,V,n$ with $n$ in unary and asks whether there is a torus tiling for $\Delta,H,V,2^{n}$, is NExpTime-complete [Für83].
We first sketch the proof idea. For a given input $\Delta,H,V,n$, we use the gadget formulae from Lemma 17 to enforce the existence of $2^{2n}$ bisimilar points $w_{0},\ldots,w_{2^{2n}-1}$. These points shall represent the $2^{n}\times 2^{n}$ cells of the exponential torus, which can be suitably addressed using propositions $b_{0},\ldots,b_{2n-1}$. Suppose for the sake of simplicity that the outdegree of the points the $w_{i}$ is at most $2^{2n}$ (instead of 3). Then we could proceed by enforcing at each $w_{i}$ with address $(x,y)$ three successors with addresses $(x,y)$, $(x,y\oplus_{2^{n}}1)$, and $(x\oplus_{2^{n}}1,y)$, and stipulate that they satisfy tile types (encoded using propositions in $\sigma$) compatible with $H,V$. Then each point in the torus is stipulated three times (from different cells), and due to the bisimilarity of the $w_{i}$ all stipulated copies of the same point satisfy the same tile type. Hence, existence of a torus tiling will coincide with some joint consistency condition. We will make this now more precise, lifting also the assumption on the outdegree.
Let $\Delta,H,V,n$ be the input to the exponential torus tiling problem. We will provide formulae $\varphi_{n},\varphi^{\prime}_{n}$ of modal depth $4n$ and with common signature $\sigma=\text{sig}(\varphi_{n})\cap\text{sig}(\varphi^{\prime}_{n})$ such that:
$\varphi_{n},\varphi^{\prime}_{n}$ are joint $\leftrightarroweq^{4n}_{\sigma}$-consistent iff there is a torus tiling for $\Delta,V,H,2^{n}$.
The common signature $\sigma$ will consist of propositions $a,b$, and one proposition $t_{d}$ for every $d\in\Delta$. Both $\varphi_{n}$ and $\varphi_{n}^{\prime}$ will use auxiliary propositions to encode counters. The formulae $\varphi_{n},\varphi_{n}^{\prime}$ are based on the families of formulae $(\psi_{i})_{i\in\mathbb{N}},(\psi_{i}^{\prime})_{i\in\mathbb{N}}$ defined in Lemma 17 and will take the shape:
$\varphi_{n}$ $=\psi_{2n}\wedge\Box^{2n}\chi_{1}$
$\varphi_{n}^{\prime}$ $=\psi_{2n}^{\prime}\wedge\Box^{2n}\chi_{2}$
for formulae $\chi_{1},\chi_{2}$ to be defined below.
Consider models $\mathcal{M}\models\psi_{2n}$ and $\mathcal{M}^{\prime}\models\psi_{2n}^{\prime}$ with $\mathcal{M}\leftrightarroweq_{\sigma}^{4n}\mathcal{M}^{\prime}$. Let $w_{0},\ldots,w_{2^{2n}-1}$ be the points in $\mathcal{M}$ and $\widehat{w}$ be the point in $\mathcal{M}^{\prime}$ that exist due to Lemma 17. Recall that, by the lemma, all $w_{i}$ are linked to $\widehat{w}$ by a $(\sigma,2n)$-bisimulation. We associate two numbers $x_{i},y_{i}$ with each point $w_{i}$ as follows:
  • $x_{i}$ is the number encoded by the valuation of $b_{0},\ldots,b_{n-1}$ in $w_{i}$, and
  • $y_{i}$ is the number encoded by the valuation of $b_{n},\ldots,b_{2n-1}$ in $w_{i}$.
Note that by the properties of $w_{0},\ldots,w_{2^{2n}-1}$, for every pair $x,y$ with $0\leq x,y<2^{n}$, there is some $w_{i}$ with $x=x_{i}$ and $y=y_{i}$. We denote that point with $w(x,y)$. Hence, the numbers $x_{i},y_{i}$ can serve as addresses of the $2^{n}\times 2^{n}$ cells in the intended torus tiling. As described above, we will exploit that all $w_{i}$ are linked to $w$ by a $(\sigma,2n)$-bisimulation to synchronize the tile types in each cell. For what follows, it is convenient to denote with $b_{0}^{i},\ldots,b_{2n-1}^{i}$ the values of propositions $b_{j}$ in point $w_{i}$, and with $c_{0}^{i},\ldots,c_{2n-1}^{i}$ and $d_{0}^{i},\ldots,d_{2n-1}^{i}$ the value of the propositions $b_{j}$ in the encoding of $x_{i},y_{i}\oplus_{2^{n}}1$ and $x_{i}\oplus_{2^{n}}1,y_{i}$, respectively. Using auxiliary non-$\sigma$ propositions, it is not difficult to write formulae $\chi_{1}$ and $\chi_{2}$ (of polynomial size) in modal logic that express the following Conditions 1 and 2, respectively. Recall that $\sigma$ contains $a,b$ and propositions $t_{d}$ for every $d\in\Delta$.
  1. (1)
    There are three paths $p_{1},p_{2},p_{3}$ of length $2n$ with the following properties:
    1. (a)
      each path satisfies $a$ in each point;
    2. (b)
      on $p_{1}$, the $j$-th point satisfies $b$ iff $b_{j-1}^{i}=1$, for $1\leq j\leq 2n$;
    3. (c)
      on $p_{2}$, the $j$-th point satisfies $b$ iff $c_{j-1}^{i}=1$, for $1\leq j\leq 2n$;
    4. (d)
      on $p_{3}$, the $j$-th point satisfies $b$ iff $d_{j-1}^{i}=1$, for $1\leq j\leq 2n$;
    5. (e)
      the ends of $p_{1},p_{2},p_{3}$ are labeled with propositions $t_{d_{1}},t_{d_{2}},t_{d_{3}}$, respectively, such that $(d_{1},d_{2})\in H$ and $(d_{1},d_{3})\in V$.
  2. (2)
    There is a (ternary) tree of depth $2n$ with the following properties:
    1. (a)
      each node has three successors: one satisfying $\neg a$, one satisfying $a,b$, and one satisfying $a,\neg b$;
    2. (b)
      in the leaves of this tree, at most one proposition $t_{d}$ is true.
This finishes the definition of the formulae $\varphi_{n},\varphi_{n}^{\prime}$ and we can proceed with showing the correctness of the reduction.
{clm}
$\psi_{2n}\wedge\Box^{2n}\chi_{1}$ und $\psi^{\prime}_{2n}\wedge\Box^{2n}\chi_{2}$ are joint $\leftrightarroweq_{\sigma}^{4n}$-consistent iff there is a torus tiling for $\Delta,H,V,2^{n}$.
To prove the claim, suppose first that $\psi_{2n}\wedge\Box^{2n}\chi_{1}$ und $\psi^{\prime}_{2n}\wedge\Box^{2n}\chi_{2}$ are joint $\leftrightarroweq_{\sigma}^{4n}$-consistent, witnessed by models $\mathcal{M},\mathcal{M}^{\prime}$. Let also be $w_{0},\ldots,w_{2^{2n}-1}$ and $\widehat{w}$ be the points that exist in $\mathcal{M},\mathcal{M}^{\prime}$ due to Lemma 17.
We define a torus tiling $\tau$ as follows. Let $x,y$ be any cell of the exponential torus, that is, $0\leq x,y<2^{n}$. Conditions 1b–1d enforce that, three paths are stipulated: one (via 1b) in point $w(x,y)$, one (via 1c) in point $w(x\ominus_{2^{n}}1,y)$, and one (via 1d) in point $w(x,y\ominus_{2^{n}}1)$, where, similar to $\oplus_{2^{n}}$, $\ominus_{2^{n}}$ denotes subtraction modulo $2^{n}$. Each of these paths is labeled, using $b\in\sigma$ with (the encoding of) $x,y$ along its elements. Due to Conditions 1a and 2a, these three paths (in $\mathcal{M}$) can only be bisimilar to the path (in $\mathcal{M}^{\prime}$) in the tree stipulated below $\widehat{w}$ that is labeled with (the encoding of) $x,y$. In particular, the ends of the paths are bisimilar to the same leaf in the tree. Since, by Condition 2b, every leaf in the tree satisfies at most one $t_{d}$, and, by Condition 1e, the end of each path satisfies at least one $t_{d}$, all ends are labeled with the same $t_{d}$. We set $\tau(x,y)=d$. Synchronization is then achieved by Condition 1e. Indeed, consider $(x,y)$ and $(x\oplus_{2^{n}}1,y)$. Then Condition 1e ensures that the end of the paths for $(x,y)$ and $(x\oplus_{2^{n}}1,y)$ stipulated in $w(x,y)$ are labeled with $t_{d},t_{d^{\prime}}$ such that $(d,d^{\prime})\in V$. The argument for horizontal compatibility is symmetric.
In the other direction, it is not difficult to construct joint $\leftrightarroweq^{4n}_{\sigma}$-consistent models from a given torus tiling for $\Delta,H,V,2^{n}$. This concludes the proof of the claim and thus of Theorem 1 as well.
We show next that the situation for the more general $\mathsf{ML}$-separability, for which the input formulae are in $\mu\mathsf{ML}$, is even worse.
{restatable}
thmthmternarycomplexity For every $d\geq 3$, $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae over $\mathbb{T}^{d}$ is 2-ExpTime-complete.
Thus, over $\mathbb{T}^{d}$ for $d\geq 3$, $\mathsf{ML}$-separability is provably harder than $\mathsf{ML}$-definability—recall that $\mathsf{ML}$-definability was shown to be ExpTime-complete in Theorem 3.1. Both the lower and the upper bound of Theorem 6.2 are non-trivial. We provide proofs for both of them in the following two subsections: first the lower bound in Subsection 6.1 and then the upper bound in Subsection 6.2. We also observe in the end of the latter Subsection 6.2 that the procedure can be easily made constructive.
6.1. Proof of Lower Bound in Theorem 6.2
We proceed in two steps. We first show a lower bound for a slightly modified separability problem, and then reduce this to $\mathsf{ML}$-separability. In detail, signature-restricted $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae is the problem that takes as an input $\mu\mathsf{ML}$-formulae $\varphi,\varphi^{\prime}$ and a signature $\sigma$ and decides whether there is an $\mathsf{ML}_{\sigma}$-separator for $\varphi,\varphi^{\prime}$. Hence, it is the variant of $\mathsf{ML}$-separability in which we can additionally restrict the signature of the sought separator.
Lemma 18.
For any $d\geq 3$, signature-restricted $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae over $\mathbb{T}^{d}$ is 2-ExpTime-hard.
Again, we show the lower bound only for $d=3$; the proof for $d>3$ is analogous. The overall proof strategy is inspired by recently studied interpolant existence problems for description logics [AJM+23], but the details have to be adapted to the present setup. The proof is via a reduction of the word problem for languages recognized by exponentially space bounded, alternating Turing machines, which we recall next.
An alternating Turing machine (ATM) is a tuple $\mathbb{A}=(Q,\Theta,\Gamma,q_{0},\Delta)$ where $Q=Q_{\exists}\uplus Q_{\forall}$ is a finite set of states partitioned into existential states $Q_{\exists}$ and universal states $Q_{\forall}$. Further, $\Theta$ is the input alphabet and $\Gamma$ is the tape alphabet that contains a blank symbol $\textsf{blank}\notin\Theta$, $q_{0}\in Q_{\forall}$ is the initial state, and $\Delta\subseteq Q\times\Gamma\times Q\times\Gamma\times\{L,R\}$ is the transition relation. We assume without loss of generality that the set $\Delta(q,a):=\{(q^{\prime},a^{\prime},M)\mid(q,a,q^{\prime},a^{\prime},M)\in\Delta\}$ contains exactly zero or two transitions for every $q\in Q_{\exists}\cup Q_{\forall}$ and $a\in\Gamma$. Moreover, the state $q^{\prime}$ in the transition $(q,a,q^{\prime},a^{\prime},M)$ must be from $Q_{\forall}$ if $q\in Q_{\exists}$ and from $Q_{\exists}$ otherwise, that is, existential and universal states alternate. A configuration of an ATM is a word $wqw^{\prime}$ with $w,w^{\prime}\in\Gamma^{*}$ and $q\in Q$. We say that $wqw^{\prime}$ is existential if $q$ is, and likewise for universal. Successor configurations are defined in the usual way. Note that every configuration has exactly zero or two successor configurations.
We define next when an ATM accepts an input. Note that according to our definition an ATM does not have accepting or rejecting states. Instead, acceptance is defined in terms of the existence infinite computations, which is slightly non-standard but more convenient for our purposes (and easily seen to be equivalent to the standard definition). A computation tree of an ATM $\mathbb{A}$ on input $w$ is a (possibly infinite) tree whose nodes are labeled with configurations of $\mathbb{A}$ such that
  • the root is labeled with the initial configuration $q_{0}w$;
  • if a node is labeled with an existential configuration $wqw^{\prime}$, then it has a single successor which is labeled with a successor configuration of $wqw^{\prime}$;
  • if a node is labeled with a universal configuration $wqw^{\prime}$, then it has two successors which are labeled with the two successor configurations of $wqw^{\prime}$.
An ATM $\mathbb{A}$ accepts an input $w$ if there is a computation tree of $\mathbb{A}$ on $w$ in which every branch is infinite. An ATM is called $f$-space bounded for some function $f:\mathbb{N}\to\mathbb{N}$ if in every configuration of every computation tree of $\mathbb{A}$ on an input of length $n$ the non-blank part of the tape is of length at most $f(n)$. It is well-known that there are $2^{n}$-space bounded ATMs which recognize a 2-ExpTime-hard language [CKS81].
For the reduction, let $\mathbb{A}=(Q,\Theta,\Gamma,q_{0},\Delta)$ be such a $2^{n}$-space bounded ATM and $w=a_{0}\ldots a_{n-1}$ be an input of length $n$. We will provide formulae $\varphi_{n},\varphi^{\prime}_{n}$ and signature $\sigma$ such that
$\varphi_{n},\varphi^{\prime}_{n}$ are joint $\leftrightarroweq^{m}_{\sigma}$-consistent for every $m\in\mathbb{N}$  iff $\mathbb{A}$ accepts $w$.
This suffices by Equivalence ($\mathsf{Base}$). Formulae $\varphi_{n},\varphi^{\prime}_{n}$ use the family of formulae $\psi_{i},\psi_{i}^{\prime}$ defined in Lemma 17, and are defined as
$\varphi_{n}$ $=\psi_{n}\wedge\Box^{n}\chi$
$\varphi_{n}^{\prime}$ $=\psi_{n}^{\prime}$
for a formula $\chi$ to be defined below. The signature $\sigma$ is given by
$\sigma=\{a,z,p_{\exists}^{1},p_{\exists}^{2},p_{\forall}\}\cup\{c_{\alpha}\mid\alpha\in\Gamma\cup(Q\times\Gamma)\}.$
Here, proposition $a$ is the one used in the gadget formulae $\psi_{n},\psi_{n}^{\prime}$, $z$ is an auxiliary proposition whose meaning becomes clear below, propositions $p_{\exists}^{*},p_{\forall}$ are responsible for marking configurations existential and universal, respectively, and finally each $c_{\alpha}$ describes a possible content of a single cell in a configuration: it is just a letter or a letter together with a state, also indicating the head position of the configuration.
To explain the main idea behind formula $\chi$ in $\varphi_{n}$, consider models $\mathcal{M},\mathcal{M}^{\prime}$ witnessing joint $\leftrightarroweq_{\sigma}^{m}$ consistency of $\varphi_{n},\varphi_{n}^{\prime}$ for some sufficiently large $m\geq n$. By Lemma 17, there are points $w_{0},\ldots,w_{2^{n}-1}$ in depth of $n$ $\mathcal{M}$ and $\widehat{w}$ in $\mathcal{M}^{\prime}$ which are linked by a $(\sigma,m-n)$-bisimulation. Additionally, they all satisfy $\chi$, due to the shape of $\varphi_{n}$. The job of $\chi$ is to enforce below each $w_{i}$ a computation tree of $\mathbb{A}$ on input $w$, as illustrated in Figure 2. Using auxiliary non-$\sigma$-propositions, it is routine to give a $\mu\mathsf{ML}$-formula which:
  • enforces the skeleton of a computation tree for $\mathbb{A}$, in which each configuration is modeled by a path of length $2^{n}$,
  • enforces that each point of the skeleton satisfies exactly one of the $\sigma$-propositions $c_{\alpha}$, and
  • enforces that universal configurations have two successors and that universal and existential configurations alternate.
At this point, it is a good moment to mention that the propositions $p_{\exists}^{1},p_{\exists}^{2}$ not only mark that the current configuration is existential, but also that they are the first or respectively second successor of a universal configuration, see beginning of the two existential configurations in Figure 2.
The challenge is to enforce that all successor configurations are consistent with $\mathbb{A}$’s transition relation $\Delta$, and we exploit here the bisimilarity of the $w_{i}$ to achieve this. The idea is that, below each $w_{i}$ we coordinate only the $(2^{n}-i)$-th position of each configuration with its successor configuration(s) in the computation tree, as depicted in Figure 2.111The figure contains some additional labeling that will be explained in more detail later. This is possible using auxiliary non-$\sigma$-propositions and due to the fact that we can access the index $i$ of $w_{i}$ via propositions $b_{0},\ldots,b_{n-1}$ mentioned in Lemma 17. Note that this provides sufficient coordination, since, due to the bisimilarity of the $w_{i}$, all the positions will be synchronized. As mentioned above, this overall strategy is similar to one employed in [AJM+23].
[…]
Figure 2. Model of a computation tree of $\mathbb{A}$, enforced by formula $\chi$.
We will now provide $\chi$ more concretely. It is a conjunction $\chi=\chi_{0}\wedge\Box^{*}\chi_{1}$, where $\Box^{*}\chi_{1}$ is an abbreviation for the $\mu\mathsf{ML}$-formula $\nu x.(\chi_{1}\wedge\Box x)$ expressing that $\chi_{1}$ holds in the entire subtree below the current point. Among others, $\chi$ uses propositions $c_{0},\ldots,c_{n-1}$ to implement an exponential counter called the C-Counter to address the cells in all configurations, and propositions $p_{\forall},p_{\exists}^{1},p_{\exists}^{2}$ to mark universal and existential configurations. Formula $\chi_{0}$ initializes the C-counter to 0 and marks the first configuration as universal (recall that $q_{0}\in Q_{\forall}$) using proposition $p_{\forall}$ as follows:
$\chi_{0} =\neg c_{0}\wedge\ldots\wedge\neg c_{n-1}\wedge p_{\forall}.$
Formula $\chi_{1}$ is a conjunction of several formulae in turn. The more standard ones are detailed in the following list. We use $(C=i)$, or similar expressions, as an abbreviation for the Boolean combination of the propositions $c_{0},\ldots,c_{n-1}$ that encodes value $i$.
  • One conjunct is responsible for incrementing the C-counter modulo $2^{n}$ along edges in models. This is entirely standard, so we refrain from detailing it.
  • The conjunct $\Diamond\top$ enforces infinite trees.
  • The following conjuncts enforce the structure of a computation tree of $\mathbb{A}$:
    $(C<2^{n}-1)\wedge p_{\forall}$ $\to\Box p_{\forall}$ (4)
    $(C<2^{n}-1)\wedge p_{\exists}^{i}$ $\to\Box p_{\exists}^{i}$  $i\in\{1,2\}$ (5)
    $(C=2^{n}-1)\wedge p_{\forall}$ $\to\Box(p_{\exists}^{1}\vee p_{\exists}^{2})$ (6)
    $(C=2^{n}-1)\wedge p_{\exists}^{i}$ $\to\Box p_{\forall}$ $i\in\{1,2\}$ (7)
    $(C=2^{n}-1)\wedge p_{\forall}$ $\to\Diamond z\wedge\Diamond\neg z$ (8)
    These implications enforce that all points which represent a configuration satisfy one of $p_{\forall},p_{\exists}^{1},p_{\exists}^{2}$ indicating the kind of configuration and, if existential, also a choice of the transition function. The symbol $z\in\Sigma$ enforces the branching. We refer the reader once again to Figure 2 for an illustration of the enforced structure. Note that $\circ$ marks the beginning of a configuration, that is, where the C-counter is 0.
  • The initial configuration on input $w=a_{0}\ldots a_{n-1}$ is enforced by
    $c_{q_{0},a_{0}}\wedge\Diamond(c_{a_{1}}\wedge\Diamond(c_{a_{2}}\wedge(\ldots\wedge\Diamond(c_{a_{n-1}}\wedge\Diamond\chi_{\textit{blank}})\ldots)),$
    where $\chi_{\textit{blank}}$ enforces label $c_{\textsf{blank}}$ until the end of the configuration. Again, this is entirely standard so we omit the details.
  • One conjunct enforces that each point is labeled with exactly one of the propositions $c_{\alpha}$, $\alpha\in\Gamma\cup(Q\times\Gamma)$.
The remaining conjuncts of $\chi_{1}$ are responsible for the coordination of the successor configurations. Recall that the goal is to coordinate in the subtree below $w_{i}$ the $2^{n}-i$-th position of each configuration with the $2^{n}-i$-th position of its successor configuration(s). As announced, we exploit the propositions $b_{0},\ldots,b_{n-1}$ from Lemma 17 for this purpose, assuming that in each $w_{i}$ the value encoded by these propositions is exactly $i$. We refer to the counter encoded by propositions $b_{0},\ldots,b_{n-1}$ as the B-counter. Analogously to the C-counter, the B-counter is incremented modulo $2^{n}$ along edges in models (enforced using another conjunct in $\chi_{1}$).
To coordinate successor configurations, we associate with $\mathbb{A}$ functions $f_{i}$, $i\in\{1,2\}$ that map the content of three consecutive cells of a configuration to the content of the middle cell in the $i$-th successor configuration (assuming an arbitrary order on the set $\Delta(q,a)$, for all $q,a$). In what follows, we ignore the corner cases that occur at the border of configurations; they can be treated in a similar way. Clearly, for each possible triple $(\alpha_{1},\alpha_{2},\alpha_{3})\in(\Gamma\cup(Q\times\Gamma))^{3}$, the $\mathsf{ML}$-formula $\varphi_{\alpha_{1},\alpha_{2},\alpha_{3}}=c_{\alpha_{1}}\wedge\Diamond(c_{\alpha_{2}}\wedge\Diamond c_{\alpha_{3}})$ is true at a point $v$ of the computation tree iff $v$ is labeled with $c_{\alpha_{1}}$, a successors $u$ of $v$ is labeled with $c_{\alpha_{2}}$, and a successors $t$ of $u$ is labeled with $c_{\alpha_{3}}$. In each configuration, we synchronize points with B-counter 0 by including for every $(\alpha_{1},\alpha_{2},\alpha_{3})$ and $i\in\{1,2\}$ the following implications:
$(B=2^{n}-1)\wedge(C<2^{n}-2)\wedge\varphi_{\alpha_{1},\alpha_{2},\alpha_{3}}\wedge p_{\forall}$ $\to\Box q^{1}_{f_{1}(\alpha_{1},\alpha_{2},\alpha_{3})}\wedge\Box q^{2}_{f_{2}(\alpha_{1},\alpha_{2},\alpha_{3})}$ (9)
$(B=2^{n}-1)\wedge(C<2^{n}-2)\wedge\varphi_{\alpha_{1},\alpha_{2},\alpha_{3}}\wedge p_{\exists}^{i}$ $\to\Box q^{i}_{f_{i}(\alpha_{1},\alpha_{2},\alpha_{3})}$ (10)
At this point, the importance of the superscript in $p_{\exists}^{\ast}$ becomes apparent: since different cells of a configuration are synchronized in trees below different $w_{k}$ the superscript makes sure that all trees rely on the same choice for existential configurations. Propositions $q^{i}_{\alpha}$ are used as markers (not in $\sigma$) and are propagated for $2^{n}$ steps, exploiting the C-counter. The superscript $i\in\{1,2\}$ determines the successor configuration that the symbol is referring to. After crossing the end of a configuration, the symbol $\alpha$ is propagated using propositions $q_{\alpha}^{\prime}$ (the superscript is not needed anymore because the branching happens at the end of the configuration, based on $z$). This is achieved by adding the following implications as conjuncts of $\chi_{1}$:
$(C<2^{n}-1)\wedge q_{\alpha}^{i}$ $\to\Box q_{\alpha}^{i}$ (11)
$(C=2^{n}-1)\wedge p_{\forall}\wedge q_{\alpha}^{1}$ $\to\Box(z\to A^{\prime}_{\alpha})$ (12)
$(C=2^{n}-1)\wedge p_{\forall}\wedge q_{\alpha}^{2}$ $\to\Box(\neg z\to q^{\prime}_{\alpha})$ (13)
$(C=2^{n}-1)\wedge p_{\exists}^{i}\wedge q_{\alpha}^{i}$ $\to\Box q^{\prime}_{\alpha}$ $i\in\{1,2\}$ (14)
$(B<2^{n}-1)\wedge q^{\prime}_{\alpha}$ $\to\Box q^{\prime}_{\alpha}$ (15)
$(B=2^{n}-1)\wedge q^{\prime}_{\alpha}$ $\to\Box q_{\alpha}$ (16)
For those $(q,a)$ with $\Delta(q,a)=\emptyset$, we add the conjunct
$\neg c_{q,a}.$
It remains to establish correctness of the reduction.
{clm}
The following conditions are equivalent:
  1. (1)
    $\mathbb{A}$ accepts $w$;
  2. (2)
    $\varphi_{n},\varphi_{n}^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{m}$-consistent for every $m\in\mathbb{N}$.
Proof 6.3.
“1 $\Rightarrow$ 2”. If $M$ accepts $w$, there is a computation tree of $\mathbb{A}$ on $w$. We construct two models $\mathcal{M}\models\varphi_{n}$ and $\mathcal{M}^{\prime}\models\varphi_{n}^{\prime}$ such that $\mathcal{M}\leftrightarroweq_{\sigma}\mathcal{M}^{\prime}$, which implies joint $\leftrightarroweq_{\sigma}^{m}$-consistency for every $m$. Let $\widehat{\mathcal{M}}$ be the infinite tree-shaped model that represents the computation tree of $\mathbb{A}$ on $w$ as described above, that is, configurations are represented by sequences of $2^{n}$ points and labeled by $p_{\forall},p_{\exists}^{1},p_{\exists}^{2}$ depending on whether the configuration is universal or existential, and in the latter case the superscript indicates which choice has been made for the existential state. Finally, the first points of the first successor configuration of a universal configuration is labeled with $z$. Observe that $\widehat{\mathcal{M}}$ interprets only the symbols in $\sigma$ as non-empty. Now, we obtain models $\mathcal{M}_{k}$, $k<2^{n}$ from $\widehat{\mathcal{M}}$ by interpreting non-$\sigma$-symbols as follows:
  • the C-counter starts at 0 at the root and counts modulo $2^{n}$ along each path starting in the root;
  • the B-counter starts at $k$ at the root and counts modulo $2^{n}$ along each path starting in the root;
  • the auxiliary concept names of the shape $q_{\alpha}^{i}$ and $q_{\alpha}^{\prime}$ are interpreted in a minimal way so as to satisfy implications (4)–(16). Note that, since these implications are essentially Horn formulae, there is a unique result.
Now, $\mathcal{M}^{\prime}$ is defined as follows:
  • start with a path of length $n$ in which each node satisfies $a$,
  • add one successor satisfying $\neg a,\neg c$ and one successor satisfying $\neg a,c$ to each node in the path, and
  • at the end $\widehat{w}$ of the path, attach a copy of $\widehat{M}$.
Next, obtain $\mathcal{M}$ from the $\mathcal{M}_{k}$ as follows:
  • Start with a full binary tree $\mathcal{M}_{0}$ of depth $n$,
  • add one successor satisfying $\neg a$ to each node in the tree,
  • interpret propositions $b_{0},\ldots,b_{n-1}$ in a way such that the B-counter values of the $2^{n}$ leaves $w_{0},\ldots,w_{2^{n}-1}$ (of the original binary tree) range from 0 to $2^{n}-1$, and
  • attach at each leaf $w_{k}$ the tree $\mathcal{M}_{k}$.
It can be verified that the reflexive, transitive, and symmetric closure of
  • all pairs $(u,v)$ for points $u$ satisfying $a$ in $\mathcal{M}$ at level $i$ and points $v$ satisfying $a$ in $\mathcal{M}^{\prime}$ at level $i$,
  • all pairs $(u,v)$ for points $u$ satisfying $\neg a$ in $\mathcal{M}$ at level $i$ and points $v$ satisfying $\neg a$ in $\mathcal{M}^{\prime}$ at level $i$, and
  • all pairs $(v,v^{\prime})$, with $v$ in $\widehat{\mathcal{M}}$ and $v^{\prime}$ a copy of $v$ in some tree $\mathcal{M}_{k}$
witnesses $\mathcal{M}\leftrightarroweq_{\sigma}\mathcal{M}^{\prime}$.
“2 $\Rightarrow$ 1”. Suppose $\varphi_{n},\varphi_{n}^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{m}$-consistent for every $m\in\mathbb{N}$. We aim to construct models $\mathcal{M},\mathcal{M}^{\prime}$ with $\mathcal{M}\models\varphi$, $\mathcal{M}^{\prime}\models\varphi^{\prime}$, and $\mathcal{M}\leftrightarroweq_{\sigma}\mathcal{M}^{\prime}$. Let us assume for a moment we have such $\mathcal{M},\mathcal{M}^{\prime}$. Since $\mathcal{M}\models\psi_{n}$ and $\mathcal{M}^{\prime}\models\psi_{n}^{\prime}$, by Lemma 17, there are pairwise $\sigma$-bisimilar points $w_{0},\ldots,w_{2^{n}-1}$ in depth $n$. That is, the trees starting at $w_{0},\ldots,w_{2^{n}-1}$ are bisimilar. By the shape of $\varphi,\varphi^{\prime}$, these trees are additionally models of $\chi_{0}\wedge\Box^{*}\chi_{1}$. It follows that in the tree below $w_{k}$, the cell contents of the $(2^{n}-k)$-th cell is coordinated, between any two consecutive configurations. Overall, all cell contents are coordinated and thus all trees below some $w_{k}$ contain a computation tree of $\mathbb{A}$ on input $w$ (which is solely represented with $\sigma$-symbols). Thus $\mathbb{A}$ accepts $w$.
It remains to construct $\mathcal{M},\mathcal{M}^{\prime}$. We use the standard technique of ‘skipping bisimulations’ exploiting that we work over models of finite outdegree. Let $\mathcal{M}_{m},\mathcal{M}_{m}^{\prime}$ be witnesses for joint $\leftrightarroweq_{\sigma}^{m}$-consistency of $\varphi_{n},\varphi_{n}^{\prime}$, for every $m\in\mathbb{N}$, and let $Z_{m}$ be a witnessing bisimulation. We can assume without loss of generality that all $\mathcal{M}_{m},\mathcal{M}_{m}^{\prime}$ are full ternary trees and that propositions outside $\text{sig}(\varphi_{n})\cup\text{sig}(\varphi_{n}^{\prime})$ are globally false. Hence, there are only finitely many possible valuations of points in $\mathcal{M}_{m},\mathcal{M}_{m}^{\prime}$.
We will construct a sequence of models $\mathcal{N}_{m},\mathcal{N}_{m}^{\prime}$, $m\in\mathbb{N}$ and $(\sigma,m)$-bisimulations $Z_{m}^{\prime}$ such that for all $m\in\mathbb{N}$ and all $i\leq m$:
  1. ($\ast$)
    $\mathcal{N}_{i}$ and $N_{m}$, $\mathcal{N}_{i}^{\prime}$ and $\mathcal{N}_{m}^{\prime}$, and $Z_{i}$ and $Z_{m}$ coincide when restricted to the first $i$ levels.
We proceed by induction. Since the roots of the $\mathcal{M}_{m},\mathcal{M}^{\prime}_{m}$ may only have one of finitely many valuations, there is a pair of valuations that occurs infinitely often in the roots. We ‘skip’ all those $\mathcal{M}_{m},\mathcal{M}_{m}^{\prime}$ which do not have the fixed valuations in their roots, ending up with a sequence that satisfies ($\ast$) for all $m\in\mathbb{N}$ and $i=0$.
In the inductive step, suppose we have constructed a sequence that satisfies ($\ast$) for every $m$ and all $i\leq i_{0}$ for some $i_{0}$. Hence, for every $j\geq i_{0}$, all points in level $i_{0}$ of $\mathcal{N}_{j},\mathcal{N}_{j}^{\prime}$ have the same valuations as in $\mathcal{N}_{i_{0}},\mathcal{N}_{i_{0}}^{\prime}$. Moreover, $uZ_{j}v$ iff $uZ_{m}v$ for all points $u,v$ in level $i_{0}$, and $j\geq i_{0}$. Since there are only finitely many points in level $i$ and each point has bounded outdegree, there are only finitely many possibilities of choosing the valuations of and bisimulations among the (finitely many points) in level $i_{0}+1$. Hence, one such choice is realized infinitely often, and we can skip all $\mathcal{N}_{m},\mathcal{N}_{m}^{\prime}$ with $m>i_{0}$ that do not realize this choice. After this modification, the sequence satisfies ($\ast$) for all $m$ and all $i\leq i_{0}+1$.
Recall that ${\mathcal{N}_{m}}_{|m}$ and ${\mathcal{N}_{m}^{\prime}}_{|m}$ denote the $m$-prefixes of $\mathcal{N}_{m}$ and $\mathcal{N}_{m}^{\prime}$, respectively, and let $Z_{m}^{\prime}$ be the restriction of $Z_{m}$ to these $m$-prefixes. It is tedious but not difficult to verify that
$\mathcal{M}=\bigcup_{m\in\mathbb{N}}{\mathcal{N}_{m}}_{|m},\quad\mathcal{M}^{\prime}=\bigcup_{m\in\mathbb{N}}{\mathcal{N}_{m}^{\prime}}_{|m},\quad\text{and}\quad Z=\bigcup_{m\in\mathbb{N}}Z^{\prime}_{m}$
satisfy $\mathcal{M}\leftrightarroweq_{\sigma}\mathcal{M}^{\prime}$, $\mathcal{M}\models\varphi_{n}$, and $\mathcal{M}^{\prime}\models\varphi_{n}^{\prime}$. The latter rely on the specific shape of $\varphi_{n},\varphi^{\prime}_{n}$ which contain only a largest fixpoint (in form of the $\Box^{*}$).
It remains to show that signature-restricted $\mathsf{ML}$-separability reduces to $\mathsf{ML}$-separability; we show it only over classes $\mathbb{T}^{d}$, for $d\geq 2$, but the reduction can easily be modified to work also over the class of all models.
Lemma 19.
For any $d\geq 2$, there is a polynomial time reduction of signature-restricted $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae over $\mathbb{T}^{d}$ to $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae over $\mathbb{T}^{d}$.
Proof 6.4.
Let $d\geq 2$, $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ and $\sigma$ be a signature. We construct $\mu\mathsf{ML}$-formulae $\widehat{\varphi},\widehat{\varphi}^{\prime}$ such that
$\varphi,\varphi^{\prime}$ are $\mathsf{ML}_{\sigma}$-separable over $\mathbb{T}^{d}$  iff $\widehat{\varphi},\widehat{\varphi}^{\prime}$ are $\mathsf{ML}$-separable over $\mathbb{T}^{d}$. (17)
Let $\sigma_{\varphi}=\text{sig}(\varphi)\setminus\sigma$ and $\sigma_{\varphi^{\prime}}=\text{sig}(\varphi^{\prime})\setminus\sigma$, respectively, be the sets of all propositions that occur in $\varphi$, respectively $\varphi^{\prime}$, but not in $\sigma$. Let $\widehat{a}$ be a proposition that does not occur in $\varphi,\varphi^{\prime}$. Then, set
$\widehat{\varphi}$ $=h(\varphi,\sigma_{\varphi})$
$\widehat{\varphi}^{\prime}$ $=h(\varphi^{\prime},\sigma_{\varphi^{\prime}})$
where $h(\psi,\tau)$ is the formula obtained from $\psi$ by
  • first inductively replacing each subformula $\Diamond\theta$ with $\Diamond\widehat{a}\wedge\Box(\widehat{a}\to\Diamond(\widehat{a}\wedge\theta))$ and each subformula $\Box\theta$ with $\Box(\widehat{a}\to\Box(\widehat{a}\to\theta))$, and
  • then replacing each proposition $c\in\tau$ with $\Diamond(\neg\widehat{a}\wedge\Diamond^{*}c)$, where $\Diamond^{*}c$ is the $\mu\mathsf{ML}$ formula that expresses that a point satisfying $c$ is reachable.
Intuitively, in the first step $h$ relativizes $\psi$ to $\widehat{a}$ and “skips” every second level, and the second step replaces non-$\sigma$ propositions by a formula that is not modally definable and not visible in the $\widehat{a}-$relativization. Clearly, the size of $h(\psi,\tau)$ is polynomial in the size of its inputs. We show correctness of the construction as stated in Equivalence (17).
For $\Leftarrow$”, suppose that $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}_{\sigma}$-separable. By Equivalence ($\mathsf{Base}$), for every $n\in\mathbb{N}$, there are models $\mathcal{M}_{n}\models\varphi,\mathcal{M}_{n}^{\prime}\models\varphi^{\prime}$ with $\mathcal{M}_{n}\leftrightarroweq^{n}_{\sigma}\mathcal{M}_{n}^{\prime}$. We construct models $\mathcal{N}_{n}\models\widehat{\varphi},\mathcal{N}_{n}^{\prime}\models\widehat{\varphi}^{\prime}$ with $\mathcal{N}_{n}\leftrightarroweq^{2n}\mathcal{N}_{n}^{\prime}$ as follows. The model $\mathcal{N}_{n}=(N_{n},v_{I},\to_{N},\mathsf{val}_{N})$ is obtained from $\mathcal{M}_{n}=(M_{n},v_{I},\to_{M},\mathsf{val}_{M})$ as follows:
  • it has universe
    $N_{n}=\{u,u^{\prime},u_{i}\mid u\in M_{n},i\in\mathbb{N}\};$
  • each $u^{\prime}$ is a successor of $u$ and each successor of $u$ (in $\mathcal{M}_{n}$) is a successor of $u^{\prime}$ (in $\mathcal{N}_{n}$);
  • for each $u$, the $u_{i}$, $i\in\mathbb{N}$ form an infinite path starting in uu; [^2] Here we need the assumption $d\geq 2$: in the constructed model $\mathcal{N}_{n}$, $u$ has at least two successors.
  • for each $u\in M_{n}$, the valuations of points in $N_{n}$ are defined as follows:
    $\mathsf{val}_{N}(u)$ $=\mathsf{val}_{M}(u)\cap\sigma$
    $\mathsf{val}_{N}(u^{\prime})$ $=\{\widehat{a}\}$
    $\mathsf{val}(u_{i})$ $=\begin{cases}\emptyset&\text{if $i\leq 2n$},\\ \mathsf{val}(u)\setminus\sigma&\text{if $i>2n$;}\end{cases}$
    Intuitively, the last equation means that the non-$\sigma$-proposition are copied from $u$ sufficiently far away in the infinite path to be invisible by any $2n$-bisimulation.
The model $\mathcal{N}_{n}^{\prime}$ is obtained analogously from $\mathcal{M}_{n}$. Using the game theoretic semantics of $\mu\mathsf{ML}$ it is not hard to show that:
{clm}
$\mathcal{N}_{n}\models h(\varphi,\sigma_{\varphi})$ and $\mathcal{N}_{n}^{\prime}\models h(\varphi,\sigma_{\varphi^{\prime}})$.
Let $Z$ be a $(\sigma,n)$-bisimulation witnessing $\mathcal{M}_{n}\leftrightarroweq_{\sigma}^{n}\mathcal{M}_{n}^{\prime}$. Based on the fact that, by construction of $\mathcal{N}_{n},\mathcal{N}_{n}^{\prime}$, no non-$\sigma$ proposition appears in the first $2n$ levels of $\mathcal{N}_{n},\mathcal{N}_{n}^{\prime}$, it is not difficult to verify that $Z^{\prime}$ defined by taking
$Z^{\prime}=Z\cup\{(u^{\prime},v^{\prime})\mid(u,v)\in Z\}\cup\{(u_{i},v_{i})\mid(u,v)\in Z,i\leq 2n\}$
witnesses $\mathcal{N}_{n}\leftrightarroweq^{2n}\mathcal{N}_{n}^{\prime}$. Hence, $\widehat{\varphi},\widehat{\varphi}^{\prime}$ are not $\mathsf{ML}$-separable.
For $\Rightarrow$”, suppose that $\widehat{\varphi},\widehat{\varphi}^{\prime}$ are not $\mathsf{ML}$-separable. Then they are not $\mathsf{ML}_{\tau}$-separable for $\tau=\text{sig}(\widehat{\varphi})\cup\text{sig}(\widehat{\varphi}^{\prime})$. By Equivalence ($\mathsf{Base}$), for every $n\in\mathbb{N}$, there are models $\mathcal{M}_{n}\models\widehat{\varphi}$, $\mathcal{M}_{n}^{\prime}\models\widehat{\varphi}^{\prime}$ with $\mathcal{M}_{n}\leftrightarroweq^{2n}_{\tau}\mathcal{M}_{n}^{\prime}$. We assume that in $\mathcal{M}_{n},\mathcal{M}^{\prime}_{n}$:
  1. (1)
    from a point not satisfying $\widehat{a}$ we never reach a point satisfying $\widehat{a}$,
  2. (2)
    no non-$\sigma$ proposition is satisfied in the first $2n$ levels, and
  3. (3)
    in even levels, every point satisfying $\widehat{a}$ has at most one successor satisfying $\widehat{a}$.
We argue that these assumptions can be made without loss of generality. For (1), recall that we relativized our formulae to $\widehat{a}$. Hence, if we modify $\mathcal{M}_{n}$ by making $\widehat{a}$ false in every subtree rooted at a point not satisfying $\widehat{a}$ it will still be a model of $\widehat{\varphi}$, and the same for $\mathcal{M}_{n}^{\prime}$. Also the bisimulation is not affected by this modification. For (2), observe that non-$\sigma$ propositions appear in $\widehat{\varphi},\widehat{\varphi}^{\prime}$ only under a $\Diamond^{*}$-operator, so we can delay them arbitrarily far. For (3), let $u$ be a point in an even level that has more than one successor satisfying $\widehat{a}$. It can be easily verified based on the structure of $\widehat{\varphi},\widehat{\varphi}^{\prime}$ that joining all successors into one (which has all the successors of the single points as successors) does not change $\mathcal{M}_{n}$ being a model of $\widehat{\varphi}$. We will from now on refer to this successor of $u$ satisfying $\widehat{a}$ with $u^{\prime}$. After doing the same modification for $\mathcal{M}_{n}^{\prime}$, we can again find a bisimulation.
We construct models $\mathcal{N}_{n},\mathcal{N}_{n}^{\prime}$ with universes $N_{n},N_{n}^{\prime}$ as follows:
  • The universe $N_{n}$ of $\mathcal{N}_{n}$ is the smallest set $N$ such that $N$ contains the root of $\mathcal{M}_{n}$, and if $u\in N$ then $N$ contains also all successors of $u^{\prime}$ (in $\mathcal{M}_{n}$) that satisfy $\widehat{a}$.
  • For $c\in\sigma$ and $u\in N_{n}$, we have $\mathcal{N}_{n},u\models c$ iff $\mathcal{M}_{n},u\models c$,
  • For $c\in\sigma_{\varphi}$ and $u\in N_{n}$, we have $\mathcal{N}_{n},u\models c$ iff $\mathcal{M}_{n},u\models\Diamond(\neg\widehat{a}\wedge\Diamond^{*}c)$,
The model $\mathcal{N}_{n}^{\prime}$ is constructed analogously from $\mathcal{M}_{n}^{\prime}$. Using the game theoretic semantics of $\mu\mathsf{ML}$ it is not hard to show that: {clm} $\mathcal{N}_{n}\models\varphi$ and $\mathcal{N}_{n}^{\prime}\models\varphi^{\prime}$. Let $Z$ be any $(2n,\tau)$-bisimulation witnessing $\mathcal{M}_{n}\leftrightarroweq^{2n}_{\tau}\mathcal{M}_{n}^{\prime}$. It is routine to verify that the restriction of $Z$ to $N_{n}\times N_{n}^{\prime}$ witnesses $\mathcal{N}_{n}\leftrightarroweq^{n}_{\sigma}\mathcal{N}_{n}^{\prime}$. Hence, $\varphi,\varphi$ are not $\mathsf{ML}_{\sigma}$-separable.
6.2. Proof of Upper Bound in Theorem 6.2
We show that over models of outdegree at most $d$, $\mathsf{ML}$-separability of fixpoint formulae can be solved in doubly exponential time. Let us start with establishing a technical but useful fact. For every language of $d$-ary trees $L\subseteq\mathbb{T}^{d}$ denote the language:
$\mathsf{bisQuot}(L)=\{\mathcal{M}\in\mathbb{T}^{d}\ |\ \text{there is $\mathcal{N}\in L$ and a functional bisimulation $Z:\mathcal{N}\stackrel{{\scriptstyle\mathsf{bis}}}{{\to}}{}\mathcal{M}$}\}$
of bisimulation quotients of trees from $L$.
Proposition 20.
For every NPTA $\mathcal{A}$, an NPTA $\mathcal{B}$ recognizing $\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ can be computed in time exponential in the size of $\mathcal{A}$.
Proof 6.5.
Fix an NPTA $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$. If necessary, we modify $\mathcal{A}$ so that no transition $S=(p_{1},...,p_{k})$ contains two occurrences of the same state. For every $\mathcal{M}\in\mathbb{T}^{d}$, we characterize existence of $d$-ary $\mathcal{N}\models\mathcal{A}$ with $\mathcal{N}\stackrel{{\scriptstyle\mathsf{bis}}}{{\to}}{}\mathcal{M}$ with the following parity game $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. The game has the set $M\times Q$ as positions. The pair $(v_{I},q_{I})$ consisting of the root $v_{I}$ of $\mathcal{M}$ and $q_{I}$ is the initial position. From a position $(v,q)$ first $\exists\text{ve}$ chooses $S\in\delta(q,\mathsf{val}(c))$ and a surjective map $h:S\to\{v_{1},...,v_{k}\}$ where $\{v_{1},...,v_{k}\}$ is the set of children of $v$. Then $\forall\text{dam}$ responds with a choice of $p\in S$ and the next round starts in position $(h(p),p)$. The game is a parity game: the ranks are inherited from $\mathcal{A}$ in the sense that the rank of $(v,q)$ equals $\mathsf{rank}(q)$. The game is designed so that:
$\text{$\exists\text{ve}$ wins $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$}\iff\mathcal{M}\in\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ (18)
for every $\mathcal{M}\in\mathbb{T}^{d}$. We prove (18) now.
We start with the implication $\Rightarrow$” from left to right. Assume a winning strategy $\zeta$ for $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. We construct $\mathcal{N}\in\mathbb{T}^{d}$ accepted by $\mathcal{A}$ together with $Z:\mathcal{N}\stackrel{{\scriptstyle\mathsf{bis}}}{{\to}}{}\mathcal{M}$. The universe $N$ consists of all finite $\zeta$-plays. There is an edge $\pi\to\pi^{\prime}$ from $\pi$ to $\pi^{\prime}$ if $\pi^{\prime}$ extends $\pi$ with one move of $\exists\text{ve}$ followed by a response of $\forall\text{dam}$. This means that the outdegree of $\pi$ equals the size of the transition $S$ chosen by $\zeta$ as a response to $\pi$. In particular, $\mathcal{N}$ is $d$-ary. Define $Z:N\to M$ such that $Z(\pi)$ is the point component $v$ from the last configuration $(v,q)$ of $\pi$. We complete the definition of $\mathcal{N}$ by putting $\mathsf{val}^{\mathcal{N}}(\pi)=\mathsf{val}^{\mathcal{M}}(Z(\pi))$ for all $\pi\in N$.
We claim that the function $Z$ is a bisimulation. The base condition $\mathsf{atom}$ follows immediately from the definition of $\mathsf{val}^{\mathcal{N}}$. To prove the $\mathsf{back}$ and $\mathsf{forth}$ conditions assume $\pi\in\mathcal{N}$. Denote $Z(\pi)=v$, let $v_{1},...,v_{k}$ be the children of $v$ and $h:S\to\{v_{1},...,v_{k}\}$ be the move chosen by $\zeta$ as a response to $\pi$.
To show the $\mathsf{forth}$ condition assume $\pi\to\pi^{\prime}$. Existence of the edge $\pi\to\pi^{\prime}$ implies that $\pi^{\prime}$ is of shape $\pi\cdot(h(p),p)$ for some $p\in S$. Let $v^{\prime}=h(p)$. Since $v\to v^{\prime}$ and $Z(\pi^{\prime})=v^{\prime}$ this completes the argument of the $\mathsf{forth}$ condition. Towards the $\mathsf{back}$ condition assume $v\to v^{\prime}$ for some $v^{\prime}$. Since $v^{\prime}$ is a child of $v$ and $h$ is surjective, there is $p\in S$ such that $h(p)=v^{\prime}$. Then $\pi^{\prime}=\pi\cdot(v^{\prime},p)$ is a $\zeta$-play such that $Z(\pi^{\prime})=v^{\prime}$ and $\pi\to\pi^{\prime}$. Hence, $\pi^{\prime}$ witnesses the $\mathsf{back}$ condition.
It remains to construct an accepting run $\rho:N\to Q$. For each $\pi\in N$ we define $\rho(\pi)$ as the state component of the last configuration in $\pi$. This $\rho$ is consistent with $\delta$. To show this assume $\pi$ with children $\pi_{1},...,\pi_{k}$. Denote $\pi_{i}=\pi\cdot(v_{i},p_{i})$ for each $i\leq k$ and let $(v,q)$ be the last configuration in $\pi$. There is $S\in\delta(q,\mathsf{val}^{\mathcal{M}}(v))$ such that $S=\{p_{1},...,p_{k}\}$. Since $\rho(\pi)=q$, $\rho(\pi_{i})=p_{i}$ for every $i\leq k$ and $\mathsf{val}^{\mathcal{N}}(\pi)=\mathsf{val}^{\mathcal{M}}(v)$, the transition $S$ is legal in $\pi$. To see that $\rho$ is accepting assume an infinite path $\pi_{1},\pi_{2},$... in $\mathcal{N}$ and for each $i$ let $(v_{i},q_{i})$ be the last configuration of $\pi_{i}$ (so in particular $\rho(\pi_{i})=q_{i}$). We need to show that the sequence $q_{1}q_{2}$... of states satisfies the parity condition. This is true because each $\pi_{i}$ is a $\zeta$-play and therefore so is their infinite limit $(v_{1},q_{1})(v_{2},q_{2})...=\pi$. This completes the proof of the implication $\Rightarrow$” in (18).
Let us prove the other implication $\Leftarrow$” in (18). Assume $\mathcal{N}\in\mathbb{T}^{d}$ such that $\mathcal{N}\models\mathcal{A}$ and $Z:\mathcal{N}\stackrel{{\scriptstyle\mathsf{bis}}}{{\to}}{}\mathcal{M}$, and let $\rho:N\to Q$ be an accepting run witnessing $\mathcal{N}\models\mathcal{A}$. We construct a winning strategy $\zeta$ for $\exists\text{ve}$ in $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. The constructed strategy preserves as an invariant that for every $\zeta$-play $\pi=(v_{1},q_{1})...(v_{l},q_{l})$ there is a path $w_{1}...w_{l}$ in $\mathcal{N}$ with $Z(w_{i})=v_{i}$ and $\rho(w_{i})=q_{i}$ for each $i\leq l$. The invariant holds in the initial position $(v_{I},q_{I})$. To define moves dictated by $\zeta$ assume a play $\pi=(v_{1},q_{1})...(v_{l},q_{l})$ and a path $w_{1}...w_{l}$ from the invariant. Let $v_{1}^{\prime},...v_{k}^{\prime}$ be the children of $v_{l}$. We define an $\exists\text{ve}$’s move $h:S\to\{v_{1}^{\prime},...,v_{k}^{\prime}\}$ dictated by $\zeta$ as a response to $\pi$.
Let $S$ be the transition chosen by $\rho$ in $w_{l}$. By assumption no state of $\mathcal{A}$ appears more than once in a single transition $S$ of $\delta$. This means that $\rho$ is bijective between the set of children of $w_{l}$ and $S$: for each $p\in S$ there exists a unique child $w^{p}$ of $w_{l}$ such that $\rho(w^{p})=p$. We set $h(p)=Z(w^{p})$ for every $p\in S$. By the $\mathsf{atom}$ condition $w_{l}$ and $v_{l}$ have the same color so to show that this $h$ is a legal move for $\exists\text{ve}$ it suffices to show that it is surjective. By the $\mathsf{back}$ condition for every child $v^{\prime}$ of $v_{l}$ there is some child $w^{\prime}$ of $w_{l}$ with $Z(w^{\prime})=v^{\prime}$. Hence, $h(\rho(w^{\prime}))=v^{\prime}$ which proves surjectivity of $h$. Moreover, the invariant is preserved: if $\pi^{\prime}$ extends $\pi$ by $\forall\text{dam}$’s response $(h(p),p)$ to $h$ then $Z(w^{p})=h(p)$ and $\rho(w^{p})=p$ and so we extend $w_{1}...w_{k}$ with $w_{k+1}=w^{p}$.
To see that the strategy $\zeta$ is winning observe that, thanks to the invariant, for every infinite $\zeta$-play $\pi=(v_{1},q_{1})(v_{2},q_{2})$... there exists an infinite path $\overline{w}=w_{1}w_{2}$... in $\mathcal{N}$ such that $Z(w_{i})=v_{i}$ and $\rho(w_{i})=q_{i}$ for all $i$. Thus, $\exists\text{ve}$ wins $\pi$ because $\rho$ is accepting. This completes the proof of (18).
Using (18) we prove Proposition 20. It suffices to construct an automaton $\mathcal{B}$ which accepts $\mathcal{M}$ iff $\exists\text{ve}$ wins $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. By positional determinacy of parity games, to check the latter it suffices to look at positional strategies for $\exists\text{ve}$ in $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. Using standard techniques we encode $\exists\text{ve}$’s positional strategies for $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$ as colorings of $\mathcal{M}$ with $\mathcal{P}(Q\times Q)$ and construct, in time exponential in $|Q|$, an automaton $\mathcal{B}^{+}$ recognizing models labelled with such winning positional strategies. We then obtain $\mathcal{B}$ recognizing $\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ by projecting out the additional colors $\mathcal{P}(Q\times Q)$ from $\mathcal{B}^{+}$.
With the help of Proposition 20 we complete the proof of Theorem 6.2. Fix $d$ and $\mu\mathsf{ML}$-formulae $\varphi$ and $\varphi^{\prime}$ over signature $\sigma$. As discussed in Section 3, it is enough to look for $\mathsf{ML}$ separators which only use symbols from $\sigma$. By Equivalence ($\mathsf{Base}$), it suffices to check if $\varphi$ and $\varphi^{\prime}$ are jointly $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for every $n$. However, unlike with definability or in the binary case, we cannot conclude joint $\cong_{\sigma}^{n}$-consistency from joint $\leftrightarroweq_{\sigma}^{n}$-consistency. Instead, we use Proposition 20 to directly decide joint $\leftrightarroweq_{\sigma}^{n}$-consistency for all $n$. For a language $L\subseteq\mathbb{T}^{d}$, define the language:
[…]
of finite $d$-ary trees which are bisimulation quotients of finite prefixes of models from $L$. By Proposition 20 and the closure properties of parity automata, for every $\mathcal{A}$ one can construct in exponential time an automaton $\mathcal{B}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$.
We prove the upper bound from Theorem 6.2. Using Theorem 1 compute automata $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$. Use Proposition 20 to compute $\mathcal{B},\mathcal{B}^{\prime}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$ and $\mathsf{QPL}(\mathcal{L}(\mathcal{A}^{\prime}))$. Recall that any two trees are bisimilar iff they have isomorphic bisimulation quotients. It follows that $\varphi,\varphi^{\prime}$ admit an $\mathsf{ML}^{n}$-separator over $\mathbb{T}^{d}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent iff $\mathcal{B},\mathcal{B}^{\prime}$ are joint $\cong^{n}_{\sigma}$ consistent. By Proposition 4, the latter condition holds for all $n\in\mathbb{N}$ iff it holds for $n=|\mathcal{B}|\times|\mathcal{B}^{\prime}|+1$ and this can be tested in time polynomial in $|\mathcal{B}|+|\mathcal{B}^{\prime}|$. Since $\mathcal{A},\mathcal{A}^{\prime}$ are exponential, and $\mathcal{B},\mathcal{B}^{\prime}$ are doubly exponential in the size of $\varphi,\varphi^{\prime}$, this gives the upper bound from Theorem 6.2.
We conclude this subsection with an observation that in case an $\mathsf{ML}$-separator exists, we can also compute one in elementary time.
{restatable}
thmthmternaryconstruction If $\varphi,\varphi^{\prime}$ are $\mathsf{ML}$-separable over $\mathbb{T}^{d}$, $d\geq 3$, then one can compute an $\mathsf{ML}$-separator in time triply exponential in $|\varphi|+|\varphi^{\prime}|$.
Proof 6.6.
It follows from the upper bound proof of Theorem 6.2 that if $\varphi,\varphi^{\prime}$ admit an $\mathsf{ML}$-separator, then they admit one of modal depth bounded doubly exponentially in $|\varphi|+|\varphi^{\prime}|$. Observe that over the signature of $\varphi$ and $\varphi^{\prime}$ there are only triple exponentially many trees of fixed outdegree $d$ and double exponential depth, and that each such tree is characterized (up to bisimulation) by a modal formula of triply exponential size. The sought separator is then the disjunction of all such formulae consistent with $\varphi$.
7. Case Study: Graded Modalities
In this section we apply our techniques and results to the case with graded modal operators. We extend $\mu\mathsf{ML}$ with formulae of the shape $\Diamond_{\geq g}\psi$ and $\Box_{\geq g}\psi$, where $g\in\mathbb{N}$ is a natural number called a grade. We denote with $\mathsf{grML}$ and $\mathsf{gr}\mu\mathsf{ML}$ the extension of $\mathsf{ML}$ and $\mu\mathsf{ML}$, respectively, with such graded modalities. Intuitively, $\Diamond_{\geq g}\psi$ is true in a point $w$ if $w$ has at least $g$ successors satisfying $\psi$ and dually, $\Box_{\geq g}\psi$ is true in $w$ if all but at most $g$ successors satisfy $\psi$ [Fin72, Ott19]. Formally, the semantics of $\mathsf{gr}\mu\mathsf{ML}$ is given by semantic games. These games extend the semantic games used to define the semantics of $\mu\mathsf{ML}$. As in the case with no grades, the game for a formula $\varphi$ and a model $\mathcal{M}$ has $M\times\mathsf{sub}(\varphi)$ as the set of positions, the winning (parity) condition is defined as in the classical case, and so are the moves for all the positions with topmost connective other than the graded modalities. In the classical game, from $(v,\Diamond\theta)$ $\exists\text{ve}$ chooses a child $v^{\prime}$ of $v$ and the next position is $(v^{\prime},\theta)$. In $(v,\Diamond_{\geq k}\theta)$, first $\exists\text{ve}$ chooses a subset $v_{1},...,v_{k}$ of size $k$ of children of $v$, then $\forall\text{dam}$ chooses one of these children $v_{i}$ and the next round starts at $(v_{i},\theta)$. Dually, in $(v,\Box_{k\geq}\theta)$ first $\exists\text{ve}$ picks a subset $v_{1},...,v_{k}$ of at most $k$ $v$’s children, then $\forall\text{dam}$ responds with a choice of some $v^{\prime}$ not in $v_{1},...,v_{k}$ and the next position is $(v^{\prime},\theta)$. Clearly, for any $d\in\mathbb{N}$, the class $\mathbb{T}^{d}$ is $\mathsf{gr}\mu\mathsf{ML}$-definable by the formula $\theta_{d}=\nu x.(\Box x\wedge\Box_{\geq d}\bot)$, which is an additional motivation to study $\mathsf{grML}$ and $\mathsf{gr}\mu\mathsf{ML}$.
We use tools and techniques from the previous sections to solve separability and definability problems in the graded setting. These questions split into two categories, depending on wether grades are allowed in the separating/defining formula or only in the input. Before we inspect these two cases in turn, let us discuss two limitations of this inquiry. First, we only look at the decision problems: deciding existence of a suitable separator/definition. The problem of efficiently constructing one when it exists is a nontrivial task, at least in the case with no grades in the separator, and goes far beyond the scope of this paper. Second, in this section we assume that the grades in formulae are represented in unary. This is admittedly a serious limitation, but is needed for an exponential translation from $\mathsf{gr}\mu\mathsf{ML}$ to automata over $\mathbb{T}^{d}$, generalizing Theorem 1. It is likely that this restriction can be lifted by working with a more succinct automata model suitable for efficiently storing grades, following the approach of [KSV02]. Adapting our proofs to this setting is not immediate, however.
7.1. With Grades in the Separator
We first investigate $\mathsf{grML}$-definability and $\mathsf{grML}$-separability of $\mathsf{gr}\mu\mathsf{ML}$-formulae. We start by introducing a graded variant of bisimulations called graded bisimulations, which characterize the expressive power of $\mathsf{grML}$ [dR00]. A relation $Z$ between models is a graded bisimulation if it satisfies ($\mathsf{atom}$) and graded variants of the ($\mathsf{back}$) and ($\mathsf{forth}$) conditions of bisimulations. The graded ($\mathsf{forth}$) condition says that if $vZw$ then for every $k\in\mathbb{N}$ and pairwise different children $v_{1},...,v_{k}$ of $v$, there are pairwise different children $w_{1},...,w_{k}$ of $w$ satisfying $v_{i}Zw_{i}$ for all $i\leq k$. The graded ($\mathsf{back}$) condition is symmetric. $Z$ is a $g$-graded bisimulation if the graded ($\mathsf{forth}$) and ($\mathsf{back}$) conditions need to be satisfied only for $k\leq g$. We denote with $\mathcal{M}\leftrightarroweq_{\mathbf{grd}}\mathcal{M}^{\prime}$ (resp., $\mathcal{M}\leftrightarroweq_{\mathbf{grd}(g)}\mathcal{M}^{\prime}$) the fact that there is a graded bisimulation (resp., a $g$-graded bisimulation) between $\mathcal{M}$ and $\mathcal{M}^{\prime}$ that relates their roots. Variants with bounded depth $n$ and/or given signature $\sigma$ are defined and denoted as expected.
{restatable}
thmthmgraded The problems of $\mathsf{grML}$-definability and $\mathsf{grML}$-separability of $\mathsf{gr}\mu\mathsf{ML}$-formulae are both ExpTime-complete.
The lower bounds follow by the usual reduction from satisfiability. We thus focus on the upper bound in the more general case of separability. Similarly to the non-graded case, we establish first a model-theoretic characterization based on graded bisimilarity.
{restatable}
lemlemgradedml For every $\varphi,\varphi^{\prime}\in\mathsf{gr}\mu\mathsf{ML}$ with maximal grade $g_{\text{max}}$, signature $\sigma$, and $n\in\mathbb{N}$, the following are equivalent:
  1. (1)
    $\varphi,\varphi^{\prime}$ are not $\mathsf{grML}^{n}_{\sigma}$-separable (over all models).
  2. (2)
    $\varphi,\varphi^{\prime}$ are joint $\leftrightarroweq^{n}_{\mathbf{grd},\sigma}$-consistent (over all models).
  3. (3)
    $\varphi,\varphi^{\prime}$ are joint $\cong^{n}_{\sigma}$-consistent (over all models).
  4. (4)
    $\varphi,\varphi^{\prime}$ are joint $\cong^{n}_{\sigma}$-consistent over $\mathbb{T}^{d}$ for $d=g_{\text{max}}\times(|\varphi|+|\varphi^{\prime}|)$.
Using Lemma 7.1 one can solve $\mathsf{grML}$-separability of $\mathsf{gr}\mu\mathsf{ML}$-formulae in exponential time, following the approach described in Section 3. More precisely, given $\varphi,\varphi^{\prime}$, we construct NPTA $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$ over $d$-ary trees, $d$ as in Lemma 7.1, and decide whether $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n$ via Proposition 4. It therefore suffices to prove the central Lemma 7.1.
Proof 7.1.
We show the implications $1\Rightarrow 2\Rightarrow 3\Rightarrow 4\Rightarrow 1$ in turn. The implication $4\Rightarrow 1$ is immediate.
For $1\Rightarrow 2$, suppose $\varphi,\varphi^{\prime}$ are not $\mathsf{grML}^{n}_{\sigma}$-separable. Hence, for every $g\in\mathbb{N}$ there is a pair of models $\mathcal{M}_{g}\models\varphi$ and $\mathcal{M}_{g}^{\prime}\models\varphi^{\prime}$ with $\mathcal{M}_{g}\leftrightarroweq_{\mathsf{grd}(g),\sigma}^{n}\mathcal{M}_{g}^{\prime}$. One can encode with an $\mathsf{FO}$-sentence $\theta$ that two models $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are depth-$n$ trees, $\mathcal{M}$ is a prefix of some $\mathcal{M}_{+}\models\varphi$ and $\mathcal{M}^{\prime}$ of some $\mathcal{M}_{+}^{\prime}\models\varphi^{\prime}$. If $Z$ is a fresh binary symbol, then it is also possible to encode with an infinite set $T$ of $\mathsf{FO}$-sentences that $Z$ is a graded bisimulation between $\mathcal{M}$ and $\mathcal{M}^{\prime}$: for every grade $g$ we take a separate $\mathsf{FO}$-sentence. Every finite fragment of $\{\theta\}\cup T$ only mentions finitely many grades and hence by assumption is satisfiable. Thus, by compactness of $\mathsf{FO}$, the entire $\{\theta\}\cup T$ is satisfiable. This gives us $\mathcal{M}\leftrightarroweq_{\mathbf{grd},\sigma}^{n}\mathcal{M}^{\prime}$ with extensions $\mathcal{M}_{+}\models\varphi$ and $\mathcal{M}_{+}^{\prime}\models\varphi^{\prime}$.
For $2\Rightarrow 3$, fix witnesses $\mathcal{M},\mathcal{M}^{\prime}$ of joint $\leftrightarroweq_{\mathbf{grd},\sigma}^{n}$-consistency. That is, $\mathcal{M}\leftrightarroweq_{\mathbf{grd},\sigma}^{n}\mathcal{M}^{\prime}$ and there are extensions $\mathcal{M}_{+},\mathcal{M}_{+}^{\prime}$ of $\mathcal{M},\mathcal{M}^{\prime}$ with $\mathcal{M}_{+}\models\varphi$ and $\mathcal{M}_{+}^{\prime}\models\varphi^{\prime}$. By the Löwenheim-Skolem property of $\mathsf{FO}$ we may assume that the $n$-prefixes of both models are at most countable. It remains to apply the known fact that countable trees $\mathcal{N}$ and $\mathcal{N}^{\prime}$ satisfy $\mathcal{N}\leftrightarroweq_{\mathbf{grd}}\mathcal{N}^{\prime}$ iff $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are isomorphic. For the sake of completeness, we add a brief justification of this latter statement. Assume $w\in\mathcal{N}$ and $w^{\prime}\in\mathcal{N}$ with respective children $\{w_{1},w_{2},...\}=W$ and $\{w_{1}^{\prime},w_{2}^{\prime},...\}=W^{\prime}$ such that $w\leftrightarroweq_{\mathbf{grd}}w^{\prime}$. For every $\leftrightarroweq_{\mathbf{grd}}$-equivalence class $X$ of $W$ the corresponding equivalence class $X^{\prime}=\{w^{\prime}\in W^{\prime}\ |\ \exists_{w\in X}.\ w\leftrightarroweq_{\mathbf{grd}}w^{\prime}\}\subseteq W^{\prime}$ has the same cardinality as $X$. This is immediate for finite $X$, and for infinite $X$ it follows because in countable models every two infinite subsets have the same cardinality. This allows us to inductively pick a bijective subrelation $Z$ of $\leftrightarroweq_{\mathbf{grd}}$ between $\mathcal{N}$ and $\mathcal{N}^{\prime}$ which is still a graded bisimulation.
For $3\Rightarrow 4$, fix witnesses $\mathcal{M},\mathcal{M}^{\prime}$ of joint $\cong_{\sigma}^{n}$-consistency. That is, $\mathcal{M}\cong_{\sigma}^{n}\mathcal{M}^{\prime}$ with respective extensions $\mathcal{M}_{+},\mathcal{M}_{+}^{\prime}$ such that $\mathcal{M}_{+}\models\varphi$ and $\mathcal{M}_{+}^{\prime}\models\varphi^{\prime}$. We trim $\mathcal{M}_{+}$ and $\mathcal{M}_{+}^{\prime}$ so that the outdegree becomes at most $d$. Without loosing generality we assume that the prefixes of $\mathcal{M}_{+}$ and $\mathcal{M}_{+}^{\prime}$ are not only isomorphic but identical.
Let $\zeta$ and $\zeta^{\prime}$ be positional winning strategies for $\exists\text{ve}$ in the semantic games $\mathcal{G}(\mathcal{M}_{+},\varphi)$ and $\mathcal{G}(\mathcal{M}_{+}^{\prime},\varphi^{\prime})$. We take a submodel $\mathcal{M}_{0}\models\varphi$ of $\mathcal{M}_{+}$ as follows. In the $n$-prefix we take the root and all $\Diamond$-witnesses for $\zeta$ and $\zeta^{\prime}$: points chosen by $\zeta$ or $\zeta^{\prime}$ in a position whose formula component starts with $\Diamond_{\geq g}$. In the rest of the model we only take $\Diamond$-witnesses for $\zeta$. A submodel $\mathcal{M}_{0}^{\prime}$ of $\mathcal{M}_{+}^{\prime}$ is defined symmetrically. It follows that $\mathcal{M}_{0}\models\varphi$ and $\mathcal{M}_{0}^{\prime}\models\varphi^{\prime}$.
Recall that $g$ is the maximal grade appearing in $\varphi$ and $\varphi^{\prime}$. Since the respective sets of positions of $\mathcal{G}(\mathcal{M}_{+},\varphi)$ and $\mathcal{G}(\mathcal{M}_{+}^{\prime},\varphi^{\prime})$ are $M_{+}\times\mathsf{sub}(\varphi)$ and $M_{+}^{\prime}\times\mathsf{sub}(\varphi^{\prime})$, for every point $v$ there are at most $g\times|\varphi|$ $\Diamond$-witnesses chosen by $\zeta$ from a position which has $v$ on the first coordinate. Consequently, the outdegree of $\mathcal{M}_{0}$ and $\mathcal{M}_{0}^{\prime}$ is not greater than $d=g\times(|\varphi|+|\varphi^{\prime}|)$. This proves Lemma 7.1.
7.2. Without Grades in the Separator
We now consider the other case with no grades allowed in the separator/definition. Interestingly, this change in the setting increases the complexity of separability, but not definability.
We start with the easier $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae, following a strategy pursued for deciding $\mathsf{ML}$-definability of $\mathsf{grML}$-formulae [KTWZ25]. As an intermediate step, and because it is interesting in its own right, we also look at $\mu\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae. It turns out that if a $\mathsf{gr}\mu\mathsf{ML}$-formula is $\mu\mathsf{ML}$-definable then this is already witnessed by the most naive candidate for a definition. Let flattening of a given $\varphi\in\mathsf{gr}\mu\mathsf{ML}$, denoted $\mathsf{flat}(\varphi)$, be a $\mu\mathsf{ML}$-formula obtained from $\varphi$ by replacing every $\Diamond_{\geq k}$ with $\Diamond$ and every $\Box_{\geq k}$ with $\Box$.
Lemma 21.
A $\mathsf{gr}\mu\mathsf{ML}$-formula $\varphi$ is equivalent to a $\mu\mathsf{ML}$-formula iff it is equivalent to $\mathsf{flat}(\varphi)$.
Proof 7.2.
We prove the only nontrivial implication from left to right. Assume $\varphi$ is equivalent to some $\mu\mathsf{ML}$-formula. In particular, this means that $\varphi$ is invariant under bisimulation. Consider a model $\mathcal{M}$. Let $\mathcal{M}^{+}$ be obtained from $\mathcal{M}$ by duplicating subtrees so that every node has infinitely many bisimilar siblings. We have:
$\mathcal{M}\models\varphi\hskip 2.84544pt$ $\iff\hskip 2.84544pt\mathcal{M}^{+}\models\varphi$
$\iff\hskip 2.84544pt\mathcal{M}^{+}\models\mathsf{flat}(\varphi)$
$\iff\hskip 2.84544pt\mathcal{M}\models\mathsf{flat}(\varphi).$
The first and third equivalence follow from bisimulation invariance. The second one follows from the observation that in $\mathcal{M}^{+}$ the semantics of $\Diamond_{\geq k}$ coincides with that of $\Diamond$, and similarly with $\Box_{\geq k}$ and $\Box$.
Lemma 21 reduces $\mu\mathsf{ML}$-definability of a given $\mathsf{gr}\mu\mathsf{ML}$-formula $\varphi$ to checking the validity of $\varphi\iff\mathsf{flat}(\varphi)$. Moreover, $\varphi$ is $\mathsf{ML}$-definable iff (i) it is equivalent to its flattening $\mathsf{flat}(\varphi)$ and (ii) $\mathsf{flat}(\varphi)$ is $\mathsf{ML}$-definable. Since the validity of $\varphi\iff\mathsf{flat}(\varphi)$ can be decided in exponential time [CGL99, KSV02] and the latter (ii) can be decided via Theorem 3.1, we obtain the following complexities.
Theorem 22.
$\mu\mathsf{ML}$-definability and $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae are ExpTime-complete.
As mentioned, without grades in the separator general separability has higher complexity than definability.
{restatable}
thmthmhalfgraded $\mathsf{ML}$-separability of $\mathsf{gr}\mu\mathsf{ML}$-formulae is 2-ExpTime-complete.
Proof 7.3.
For the lower bound, we reduce $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae over $\mathbb{T}^{3}$. Since the former problem is 2-ExpTime-hard by Theorem 6.2, so is the latter. The reduction relies on the formula $\theta_{3}$ defining $\mathbb{T}^{3}$. It is an immediate consequence of the definitions that for all $\mu\mathsf{ML}$-formulae $\varphi,\varphi^{\prime}$ and $\psi\in\mathsf{ML}$,
[…]
Towards the upper bound, assume $\varphi,\varphi^{\prime}\in\mathsf{gr}\mu\mathsf{ML}$. As usual, we only need to look for separators among formulae over $\sigma=\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. By an immediate generalization of the Equivalence ($\mathsf{Base}$), $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}_{\sigma}^{n}$-separable iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent. A standard game-theoretic argument similar to the one in the proof of $3\Rightarrow 4$ in Lemma 7.1 shows that $\varphi,\varphi$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over all models iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$, where $d=g\times(|\varphi|+|\varphi^{\prime}|)$ and $g$ is the greatest grade occurring in $\varphi,\varphi^{\prime}$. Summing up, $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}$-separable iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n\in\mathbb{N}$. To decide the latter we proceed as in the upper bound proof of Theorem 6.2. We first construct NPTA $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$ over $d$-ary trees via (an analogue for $\mathsf{gr}\mu\mathsf{ML}$ of) Theorem 1. Then, using Proposition 20 we construct $\mathcal{B}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$ and $\mathcal{B}^{\prime}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}^{\prime}))$ and apply Proposition 4 to decide if $\mathcal{B},\mathcal{B}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent for all $n$.
8. Final Discussion
We have presented an in-depth study of modal separation of $\mu\mathsf{ML}$-formulae over different classes of structures. For us, the most interesting results are the differences that are obtained over classes of bounded outdegree for different bounds $d=1$, $d=2$, $d\geq 3$. It is worth noting that without much effort our (upper bound) results can be transferred to other interesting classes of models. First, it is known that $\mu\mathsf{ML}$ has the tree model property and the finite model property. Due to this, for every $\psi,\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ it follows that $\psi$ is a separator of $\varphi,\varphi^{\prime}$ (over all models) iff $\psi$ is a separator of $\varphi,\varphi^{\prime}$ over trees iff $\psi$ is a separator of $\varphi,\varphi^{\prime}$ over finite models. Thus, $\mathsf{ML}$-separability coincides over all these classes. Moreover, with the help of the $\mu\mathsf{ML}$-formula $\theta_{\infty}$ from Example 3 we can show that $\mathsf{ML}$-separability over finite trees reduces to $\mathsf{ML}$-separability over all models. More formally:
Lemma 23.
Let $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ and $\psi\in\mathsf{ML}$. Then $\psi$ is an $\mathsf{ML}$-separator of $\varphi,\varphi^{\prime}$ over finite trees iff $\psi$ is an $\mathsf{ML}$-separator of $\varphi\wedge\neg\theta_{\infty},\varphi^{\prime}\wedge\neg\theta_{\infty}$. This is also true inside $\mathbb{T}^{d}$, for $d\in\mathbb{N}$.
Via similar reductions, our upper bound results do also transfer to infinite words and ranked trees.
We find also worth noting that the construction of $\mathsf{ML}^{n}$-uniform consequences of $\mu\mathsf{ML}$ formulae might be of independent interest. For example, the construction can easily be adapted to construct $\mathsf{ML}^{n}_{\sigma}$-uniform consequences, which in turn provides an alternative route to computing (uniform) interpolants of $\mathsf{ML}$-formulae.
We discuss a few more aspects and possible extensions of our work.
8.1. Craig Separability.
Inspired by Craig interpolation, it is natural to also consider the following variant of separability. Given $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$, we say that a formula $\psi\in\mathsf{ML}$ is a Craig $\mathsf{ML}$-separator of $\varphi,\varphi^{\prime}$ if $\text{sig}(\psi)\subseteq\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$ and $\psi$ separates $\varphi,\varphi^{\prime}$. Craig $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae is then the problem of deciding whether given $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ admit a Craig $\mathsf{ML}$-separator. Interestingly, $\mathsf{ML}$-separability and Craig $\mathsf{ML}$-separability coincide over all models, over $\mathbb{T}^{1}$, and over $\mathbb{T}^{2}$.
Proposition 24.
Over the class of all models, as well as over $\mathbb{T}^{1}$ and $\mathbb{T}^{2}$, $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ admit an $\mathsf{ML}$-separator iff they admit a Craig $\mathsf{ML}$-separator.
Proof 8.1.
The “if”-direction is trivial. Clearly, any Craig $\mathsf{ML}$-separator is also an $\mathsf{ML}$-separator.
For “only if” we will use the fact that $\mathsf{ML}$ enjoys the CIP over each of the considered classes $\mathbf{C}$: over all models and over $\mathbb{T}^{1}$ [Gab72], and over $\mathbb{T}^{2}$ by Proposition 12. Suppose there is an $\mathsf{ML}$-separator $\psi$ of $\varphi,\varphi^{\prime}$ and let $n$ be its modal depth. Let $\theta,\theta^{\prime}$ be respective $\mathsf{ML}^{n}$-uniform consequences of $\varphi,\varphi^{\prime}$, which exist by Proposition 7 or Proposition 10 depending on the class $\mathbf{C}$. Note that in both cases we have $\text{sig}(\theta)\subseteq\text{sig}(\varphi)$ and $\text{sig}(\theta^{\prime})\subseteq\text{sig}(\varphi^{\prime})$. Moreover, since $\psi\in\mathsf{ML}^{n}$, we have $\varphi\models\theta\models\psi\models\neg\theta^{\prime}\models\neg\varphi^{\prime}$. Thus, a Craig interpolant for $\theta,\neg\theta^{\prime}$ is a Craig separator for $\varphi,\varphi^{\prime}$.
Clearly, Proposition 24 cannot hold over ternary (and higher arity) trees, since $\mathsf{ML}$ does not even enjoy the CIP over $\mathbb{T}^{3}$ for $d\geq 3$. It is also interesting to note that the CIP alone is not sufficient for Proposition 24: there are classes $\mathbf{C}$ where $\mathsf{ML}$ enjoys the CIP but $\mathsf{ML}$-separability and Craig $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae do not coincide. Fix a proposition $a\in\mathsf{Prop}$ and consider:
[…]
The formula $\theta_{\infty}$ is equivalent to $a$ over $\mathbf{C}$. However, $\text{sig}(\theta_{\infty})=\emptyset$ and no modal formula over the empty signature separates $\theta_{\infty},\neg\theta_{\infty}$ over $\mathbf{C}$. Hence, $\mathsf{ML}$-separability and Craig $\mathsf{ML}$-separability do not coincide over $\mathbf{C}$. On the other hand, it is not difficult to derive the CIP of $\mathsf{ML}$ over $\mathbf{C}$ from the CIP of $\mathsf{ML}$ over $\mathbb{T}^{1}$.
8.2. Other Size Measures.
Throughout the paper we used the simplest possible measure of formula size: the length of a formula written as a string. Alternative more succinct measures, such as the number of non-isomorphic subformulae (DAG-size), are also interesting. Thus, a natural question is to what extent our results depend on the choice of size measure. In principle, using a more succinct measure makes the problems of definability and separability harder. However, all our decision procedures, with an exception of Theorem 1, are automata-based. Consequently, these procedures carry over with unchanged complexity to any size measure for which the translation from logic to nondeterministic automata has the same complexity as in Theorem 1. In the remaining case of Theorem 1 a sufficient assumption is that the modal depth of a formula is at most polynomial in its size. Both the mentioned assumptions are arguably modest.
A place where the choice of size measure matters a little more is the construction of modal definitions and separators. In the cases of unrestricted, unary ($\mathbb{T}^{1}$), and high outdegree models ($\mathbb{T}^{d}$ for $d\geq 3$) the constructed formulae have DAG-size essentially the same as size: doubly, singly, and triply exponential, respectively. Interestingly, however, in the binary case $\mathbb{T}^{2}$ our formulae have only singly exponential DAG-size, which is easily seen to be optimal and contrasts with their doubly exponential size. This demonstrates that the lower bounds for size of modal definitions over $\mathbb{T}^{2}$ cannot work for DAG-size. The same lower bound construction fails for DAG-size over unrestricted models, although there the exact DAG-size complexity of optimal separators remains unknown.
8.3. Open Problems.
We mention some interesting open problems. First, the relative succinctness of $\mu\mathsf{ML}$ over $\mathsf{ML}$ is to the best of our knowledge open in the setting with only one accessibility relation. Second, as we have mentioned in Section 3, the separators we compute are not necessarily the logically strongest ones. The logically strongest separators of $\varphi,\varphi^{\prime}$ are precisely the $\mathsf{ML}$-uniform consequences of $\varphi$ (if they exist) and are a natural object of study. Clearly, modal definability of $\varphi$ is a sufficient condition, but not a necessary one. Let $\varphi=\psi\wedge\neg\theta_{\infty}$ and $\varphi^{\prime}=\neg\psi$ for some $\psi\in\mathsf{ML}$ and $\theta_{\infty}$ from Example 3. Then $\varphi$ is not equivalent to a modal formula, but $\psi$ is a strongest separator. It would also be interesting to investigate separability of $\mu\mathsf{ML}$-formulae by other fragments of $\mu\mathsf{ML}$, such as safety formulae, for which definability has been studied in [LQ15] and to extend the $\mu$-calculus by other operators, e.g., nominals or inverse modalities.
Finally, it would be interesting to extend our findings in the case study on the graded $\mu\mathsf{ML}$ in Section 7. As already pointed out, there are two natural directions for this: first one could study the complexity under binary coding of grades, and second one could tackle the separator construction problem. We conjecture that the first problem can be solved by combining our techniques with the automata constructions from [KSV02], which lead to optimal complexity for satisfiability of $\mathsf{grML}$ under binary coding of grades. The computation problem might be much more challenging. Indeed, it was observed recently that the construction of $\mathsf{ML}$ separators of $\mathsf{grML}$-formulae already requires significantly new ideas [JKW25].